### v1

# On the Effectiveness of Methods and Metrics for Explainable AI in Remote Sensing Image Scene Classification

Abstract
The development of explainable artificial intelligence (xAI) methods for scene classification problems has attracted great attention in remote sensing (RS). Most xAI methods and the related evaluation metrics in RS are initially developed for natural images considered in computer vision (CV), and their direct usage in RS may not be suitable. To address this issue, in this paper, we investigate the effectiveness of explanation methods and metrics in the context of RS image scene classification. In detail, we methodologically and experimentally analyze ten explanation metrics spanning five categories (faithfulness, robustness, localization, complexity, randomization), applied to five established feature attribution methods (Occlusion, LIME, GradCAM, LRP, and DeepLIFT) across three RS datasets. Our methodological analysis identifies key limitations in both explanation methods and metrics. The performance of perturbation-based methods, such as Occlusion and LIME, heavily depends on perturbation baselines and spatial characteristics of RS scenes. Gradient-based approaches like GradCAM struggle when multiple labels are present in the same image, while some relevance propagation methods (LRP) can distribute relevance disproportionately relative to the spatial extent of classes. Analogously, we find limitations in evaluation metrics. Faithfulness metrics share the same problems as perturbation-based methods. Localization metrics and complexity metrics are unreliable for classes with a large spatial extent. In contrast, robustness metrics and randomization metrics consistently exhibit greater stability. Our experimental results support these methodological findings.
Based on our analysis, we provide guidelines for selecting explanation methods, metrics, and hyperparameters in the context of RS image scene classification. The code of this work will be publicly available at https://git.tu-berlin.de/rsim/xai4rs.
Index Terms:
Explainable Artificial Intelligence, Remote Sensing, Deep Learning, Feature Attribution, Explanation Metrics.
I Introduction
The development of accurate methods for remote sensing (RS) image scene classification is one of the most important research topics in RS. In recent years, deep learning (DL)has emerged as the dominant approach for scene classification problems due to the high capability of DL models to extract and exploit the complex spatial and spectral content of RS images. RS image scenes can contain multiple land-use land-cover (LULC) classes and thus can be simultaneously associated with different class labels (i.e. multi-labels). Accordingly, the existing models are in general capable of assigning multiple LULC class labels to each RS image scene in an archive [1, 2]. Despite their success, DL models often remain enigmatic black boxes, with almost no human understanding of their internal processes [3]. This lack of transparency raises concerns as decision-makers increasingly rely on these black box systems [4]. New regulations for critical applications, such as the General Data Protection Regulation [5] and the Artificial Intelligence Act [6], require interpretability to ensure transparent decisions. Beyond regulatory concerns, effective interpretability can also improve model performance by addressing issues such as limited or biased training data, outliers, vulnerability to adversarial attacks, and overfitting [4].
To this end, explainable artificial intelligence (xAI)has emerged as a research area focused on revealing the factors behind DL model predictions. Although xAI methods are widely used in the CV community [7], they have only recently gained attention in RS. For a comprehensive overview of xAI in RS we refer the reader to [8]. The most prominent type of explanation methods for RS scene classification problems are feature attribution methods [8]. These methods assign importance scores to input features based on their contribution to the prediction of a model[9, 10, 11, 12, 13]. While some feature attribution methods have been adapted to RS scene classification[14, 15], most studies use feature attribution methods that were originally designed for image classification tasks on natural images [16, 17, 18].
This raises the question of whether they are applicable to RS image scenes, as they exhibit different characteristics compared to natural images and can even be considered a distinct modality [19]. To answer this question, it is essential to assess the quality of the generated explanations. Existing evaluation methods can be categorized into two groups: 1) qualitative assessments, which include domain expert evaluations or user studies; and 2) quantitative assessments, which employ explanation metrics [20]. Recently, Höhl et al. [8] reviewed approximately 200 RS publications, finding that the vast majority evaluated qualitative, while only 16 studies included quantitative evaluation using explanation metrics. This reliance on qualitative assessment raises concerns, as studies from the CV community have shown that qualitative assessment is insufficient for robust verification of explanation methods [21, 22].
One of the key reasons for the limited use of quantitative evaluations in RS is the absence of clear guidelines for assessing explanation quality in this domain. The need for such guidelines is underscored by the inherent difficulty of quantitative evaluation since explanation metrics depend heavily on their parametrization [23, 22]. Moreover, the lack of ground truth for the explanations renders them unverifiable [24], requiring the use of a wide range of metrics for a more comprehensive evaluation [25]. In many RS studies only a small part of this range is analyzed, often focusing on single aspects, such as localization capability [17], while only a few studies have evaluated a broader variety of metrics [26, 27]. In particular, none of these studies have addressed the fact that most explanation metrics have been originally designed for natural images and can be limited when applied to RS scene classification.
To address these issues, we present the first unified evaluation of the explanation methods and metrics for RS image scene classification. We assess five widely used feature attribution methods: Occlusion [9], LIME [10], LRP [11], DeepLIFT [12] and GradCAM [13]; and evaluate ten explanation metrics spanning five categories: faithfulness (FE [28], IROF [29]), robustness (AS [30], LLE [28]), localization (RRA [31], TKI [32]), complexity (SP [33], CO [34]) and randomization-based (MPRT [35], RL [36]). These methods and metrics are selected for their popularity in RS and to cover a broad spectrum of evaluation criteria. We conduct a methodological analysis to examine whether existing explanation methods and metrics are suitable for RS image scene classification. Then, we perform an extensive experimental evaluation using MetaQuantus [24] to assess the reliability and effectiveness of explanation metrics under minor perturbations and disruptive changes, supporting our methodological findings with experimentally validation.
We would like to note that, unlike previous work that primarily surveys the state of the art in xAI for RS [8], our study goes beyond a methodological overview by conducting a systematic, in-depth analysis of explanation methods and evaluation metrics. In particular, we critically examine the reliability of explanation metrics, which is an aspect that, to the best of our knowledge, has not previously been investigated in the RS domain. The primary contributions of this work are summarized as follows:
  • We provide an extensive methodological analysis of how the differences between RS and natural images impact the effectiveness of existing feature attribution methods and evaluation metrics in the context of scene classification in RS.
  • Based on this analysis, we conduct an experimental study of the selected attribution metrics and methods across multiple RS scene classification datasets.
  • Based on our methodological and experimental findings, we develop practical guidelines to assist RS practitioners in selecting appropriate explanation methods and evaluation metrics for RS image scene classification and recommend suitable explanation methods and metrics. Our pipeline can be reproduced and adapted for different datasets or classification tasks in RS, providing a valuable resource for the community.
The remainder of this paper is organized as follows. Section II and Section III provide a methodological evaluation of the effectiveness of explanation methods and explanation metrics for RS image scene classification, respectively. In Section IV we describe the datasets and experimental setup. The results of our experiments are presented in Section V. Finally, Section VI draws the conclusion of the work.
Figure 1: The evaluation protocol for selecting appropriate explanation methods and metrics for RS image scene classification. It involves generating explanations using various feature attribution methods, estimating their quality using selected metrics, assessing the reliability of the metric with MetaQuantus, and using the most reliable metrics to select the most suitable explanation methods.
II Explanation methods for RS image scene classification
In this section, we methodologically analyze the explanation methods to address the unique challenges posed by the RS images. Most xAI research has focused on natural images, which differ significantly from RS images [19]. To this end, we first analyze a range of explanation methods commonly applied in RS scene classification regarding these differences. We focus on local post-hoc explanation methods that generate feature attribution maps, as they are most commonly used in RS [8]. Feature attribution methods can be further grouped into backpropagation-based and perturbation-based methods [37]. Perturbation-based methods alter model inputs to observe changes in the model output, while backpropagation methods propagate values from the model output to the input space, typically gradients or relevancy scores. We consider three categories: i) perturbation-based methods, ii) relevancy-propagation-based methods and iii) gradient-propagation-based methods. To provide a comprehensive overview, our analysis includes five popular feature attribution methods covering all categories: 1) two perturbation-based methods (Occlusion [9], LIME [10]); 2) two relevancy-propagation-based methods (LRP [11], DeepLIFT [12]) and 3) one gradient-propagation-based method (GradCAM [13]). For each category, we first provide a general description, followed by a definition of the explanation methods and a discussion of the limitations considering the differences between RS and natural images.
Let us consider a RS image scene classification problem where a DL model $f$ is trained on a RS image archive $\bm{X}_{\mathrm{tr}}$ made up of $N$ pairs: $(\bm{x}_{i},y_{i})_{i=1}^{N}$ for single‐label classification, or $(\bm{x}_{i},\bm{y}_{i})_{i=1}^{N}$ for multi‐label classification. For simplicity, we omit the single-label notation. Each image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ represents a scene captured by an airborne or satellite sensor, where $C$ denotes the spectral bands and $H,W$ correspond to spatial dimensions. The model $f$ maps an input image $\bm{x}$ to a predicted set of labels:
$f(\bm{x})=\hat{\bm{y}}$ (1)
where $\hat{\bm{y}}$ represents the predicted class probabilities or binary assignments. To interpret the reasoning behind the prediction of a model $f$ for an input $\bm{x}$, post-hoc explanation methods, such as feature attribution methods $\Phi$, can be applied:
$\Phi(\bm{x},f,\hat{\bm{y}})=\hat{\bm{e}},$ (2)
where $\Phi$ distributes relevancy scores $r$ across $\bm{x}$, resulting in an attribution map $\hat{\bm{e}}\in\mathbb{R}^{H\times W\times C}$. In practice, $\hat{\bm{e}}$ is often aggregated across spectral bands. For simplicity, we omit explicit parameterization unless necessary.
II A Perturbation-based methods
II A 1 Background
Perturbation-based explanation methods explain model predictions by measuring changes in the output between the original and perturbed inputs. Let $\mathbb{X}\subseteq\mathbb{R}^{H\times W\times C}$ be the space of valid inputs (e.g., images) and let $\Theta$ be a set of perturbation parameters (e.g., masks, noise levels, patch locations). A perturbation function $\mathcal{P}_{\mathbb{X}}$ is defined as:
$\mathcal{P}_{\mathbb{X}}:\mathbb{X}\times\Theta\;\to\;\mathbb{X},$ (3)
such that for any $\bm{x}\in\mathbb{X}$ and any $\theta\in\Theta$, the output
$\tilde{\bm{x}}\;=\;\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$ (4)
remains in $\mathbb{X}$. In other words, $\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$ modifies $\bm{x}$ according to the perturbation parameter $\theta$, while still preserving the overall structure such that $\tilde{\bm{x}}$ is a valid point in the same space. We define the category of perturbation-based methods $\Phi_{\text{Perturbation}}$ as:
$\Phi_{\text{Perturbation}}(f,\bm{x},\mathcal{P}_{\mathbb{X}},\theta)=|f(\bm{x})-f(\mathcal{P}_{\mathbb{X}}(\bm{x},\theta))|.$ (5)
Common perturbation functions are blurring [38], introducing noise [30], or replacing parts of the image [39]. For perturbation-based explanation methods, we consider Occlusion [9] and LIME [10].
The Occlusion method [9] is defined as follows. Let each $\theta\in\Theta$ specify a region (mask) $\Omega\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ and a constant perturbation baseline $\alpha\in\mathbb{R}$. We define the perturbation function:
$\bigl{[}\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)\bigr{]}_{j,k}\;=\;\begin{cases}\alpha,&\text{if }(j,k)\in\Omega,\\ \bm{x}_{j,k},&\text{otherwise}.\end{cases}$ (6)
Then, we calculate the prediction certainty difference $\Delta s_{c}$ as:
$\Delta s_{c}(\theta)=f_{c}(\bm{x})-f_{c}\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)\bigr{)},$ (7)
for each mask $\Omega$. A large positive $\Delta s_{c}(\theta)$ means that the perturbation significantly lowered the prediction certainty of the model for class $c$, suggesting that the perturbed features are important. Then, we aggregate the $\Delta s_{c}(\theta)$ values to produce an attribution map that assigns high importance to the regions whose occlusion most reduces the model prediction certainty for class $c$. Thus, the Occlusion method $\Phi_{\text{Occlusion}}$ is defined as:
$\Phi_{\text{Occlusion}}(j,k)\;=\;\frac{1}{\bigl{|}\{\theta:(j,k)\in\Omega\}\bigr{|}}\sum_{\theta:(j,k)\in\Omega}\Delta s_{c}(\theta).$ (8)
The LIME $\Phi_{\mathrm{LIME}}$ method explains a black-box model $f$ by locally approximating it with a simple, interpretable model trained on binary-encoded perturbations of the original input. Let $S=\{S_{1},S_{2},\dots,S_{M}\}$ be a segmentation of $\bm{x}$ into $M$ superpixels. Denote by $\bm{z}\in\{0,1\}^{M}$ a binary vector that indicates which superpixels are present (1) or not present (0). We define the perturbation function:
$\mathcal{P}_{\mathbb{X}}:\mathbb{X}\times\{0,1\}^{M}\;\to\;\mathbb{X},$ (9)
such that, for each $\theta=\bm{z}$, where:
$\mathcal{P}_{\mathbb{X}}(\bm{x},\bm{z})_{j,k}\;=\;\begin{cases}[\bm{x}]_{j,k},&\text{if }(j,k)\in S_{m}\text{ ,}\bm{z}_{m}=1,\\ \alpha,&\text{if }(j,k)\in S_{m}\text{ , }\bm{z}_{m}=0,\end{cases}$ (10)
and $\alpha\in\mathbb{R}^{C}$ is a constant perturbation baseline. By toggling each superpixel via $\bm{z}_{m}\in\{0,1\}$, we create perturbed samples in the neighborhood of $\bm{x}$. LIME estimates a simple, interpretable model $g$ (e.g., a sparse linear model) that approximates $f$ on these perturbations near $\bm{x}$. Formally, let:
$\Phi_{\mathrm{LIME}}(\bm{x})\;=\;\operatorname{argmin}_{g\in\mathcal{G}}\Bigl{[}\mathcal{L}\bigl{(}f,g,\Pi_{x}\bigr{)}\;+\;\Upsilon(g)\Bigr{]}.$ (11)
Here, $\Upsilon(g)$ is a regularization term, such as the number of nonzero weights in a linear model or the depth of a decision tree, $\mathcal{L}$ is a loss function, and $\Pi_{x}(\tilde{\bm{x}})$ denotes a proximity measure that quantifies the distance from a perturbation $\tilde{\bm{x}}$ to $\bm{x}$.
II A 2 Limitations
A key component of perturbation-based methods is the perturbation baseline, which is intended to represent the absence of a feature. However, defining an appropriate perturbation baseline is non-trivial. Many RS studies simply employ a scalar value $\alpha$, often set to zero, such as in the quantitative evaluation of Kakageorgiou et al. [26]. However, perturbations with a scalar perturbation baseline can yield Out-of-Distribution (OoD)samples [40]. This is problematic because DL models are prone to unreliable predictions when the input does not reside in their learned distribution [41]. An example case is shown in Fig. 2(a), which depicts an RGB image from the DeepGlobe Multi-Label (DG-ML)dataset. A rectangular patch is perturbed with the Occlusion method, using a black perturbation baseline ($\alpha=0$). Despite occluding a part of the class, the prediction of the model for the class agricultural land increased from $f_{c}(\bm{x})=0.90$ on the original input to $f_{c}(\tilde{\bm{x}})=0.99$ after the occlusion. Even when most of the pixels from the original image are perturbed, as shown in Fig. 2(b), the model predicts with high certainty $f_{c}(\tilde{\bm{x}})=0.38$, highlighting the unpredictable behavior introduced by a poorly chosen perturbation baseline.
This challenge is amplified in multispectral RS images, where different spectral bands (e.g., near-infrared or thermal) capture distinct information. Common baselines, such as all-zero vectors, can be unsuitable because certain bands rarely contain zero values in operational conditions. For example, in thermal imagery, a zero-valued perturbation baseline is not meaningful, as it does not reflect any typical temperature reading. Similarly, for a near-infrared band, some level of infrared radiation is almost always present, making an all-zero replacement an inadequate representation of the absence of a feature. Consequently, selecting a meaningful perturbation baseline requires more sophisticated approaches that account for the different spectral bands.
Figure 2: Visualizations of perturbed images using $\alpha=0$ from the DG-ML dataset for the target class $c$ (agricultural land) and its prediction certainty for the original image: $f_{c}(\bm{{x}})=0.9$. a) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.99$: b) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.38$.
Furthermore, the regions within each RS image are associated with different LULC classes, and unlike in natural images, background content does not exist. Let us consider the case where we have a LULC class $c$ present in the image and let $A_{c}(\bm{x})$ represent the total number of pixels in $\bm{x}$ corresponding to class $c$ (which can be interpreted as the discrete spatial extent of class $c$ within the image). If the area covered by the perturbation mask $A(\Omega)$ of the Occlusion method is smaller than $A_{c}(\bm{x})$, the perturbation may fail to obscure a sufficient number of relevant pixels to significantly decrease the prediction certainty. This is depicted in Fig. 2(a), where we see a rectangular region that is occluded, which does not reduce $f_{c}(\bm{\tilde{x}})$. This limitation is less pronounced in natural images, where discriminative features are often unique (e.g., silhouette, object parts). In such cases, occluding typically leads to a noticeable reduction in prediction certainty even if $A_{c}(\bm{x})>A(\Omega)$.
A similar problem occurs for the LIME method, where $\mathcal{P}_{\mathbb{X}}$ partitions $\bm{x}$ in different superpixels and partially occludes them. Here, the prediction remains unchanged, unless most of the superpixels are replaced with the perturbation baseline (see Fig 2(b)), resulting in a linear model trained primarily on positive class predictions. This imbalance causes the estimated weights of the linear model to be unreliable, leading to an inaccurate interpretation of feature importance. Since no single superpixel significantly alters the decision of the model, LIME assigns almost equal importance to all superpixels. However, this relevance is lower than it should be, as the method underestimates the contribution of discriminative features due to the regularization term, potentially leading to wrong interpretations of the reasoning of the model. Additionally, the more pixels are replaced by the perturbation baseline, the higher the proximity measure $\Pi_{x}$ due to the large distance from the original image. As a result, individual pixels are incorrectly attributed as less relevant.
II B Gradient-propagation-based methods
II B 1 Background
Gradient-propagation-based methods utilize gradient information to generate visual explanations for deep neural network predictions. The gradient-propagation-based explanation method we examine is GradCAM [13], which incorporates gradient information flowing into any convolutional layer, typically the last one due to its high-level semantics and spatial information retention. The attribution map for a specific class $c$ is computed by first capturing the gradients of the class score $Y^{c}$ with respect to the activations $A^{p}$ of convolutional layer $l$. The importance weights $\omega^{c}_{p}$ for each feature map are determined by globally averaging the gradients over all spatial locations $(j,k)$:
$\omega^{c}_{p}=\frac{1}{Z}\sum_{j,k}\frac{\partial Y^{c}}{\partial A^{p}_{j,k}}.$ (12)
These weights represent a partial linearization of the network and denote the importance of each feature map $p$ for the target class $c$. A weighted combination of these maps followed by a ReLU ensures that only positive influences on the class prediction are visualized:
$\Phi_{\mathrm{GradCAM}}^{c}=\text{ReLU}\left(\sum_{p}\omega^{c}_{p}A^{p}\right).$ (13)
II B 2 Limitations
Gradient-propagation-based methods exhibit limitations when applied to RS imagery, particularly when multiple LULC classes occur in a scene. For instance, in a RS image containing five airplanes, GradCAM localizes only a single instance [42], underscoring its inability to resolve multiple objects of interest at varying scales [43, 44]. This failure originates from the reliance of GradCAM on a global gradient averaging mechanism, which collapses spatial gradient information into a single channel-wise importance weight. For two distinct objects $O_{1}$ and $O_{2}$, activating disjoint regions $R_{1}$ and $R_{2}$ in $A^{p}$, the gradients $\nabla_{1}$ and $\nabla_{2}$ are:
$\frac{\partial Y^{c}}{\partial A_{j,k}^{p}}=\begin{cases}\nabla_{1},&\forall(j,k)\in R_{1}\\ \nabla_{2},&\forall(j,k)\in R_{2}\\ 0,&\text{otherwise}.\end{cases}$ (14)
Assuming equal spatial extents ($\lvert R_{1}\rvert=\lvert R_{2}\rvert=N$) and total positions $Z=u\times v$, GradCAM computes the channel weight as:
$\omega_{p}^{c}=\frac{N}{Z}\,\bigl{(}\nabla_{1}+\nabla_{2}\bigr{)}.$ (15)
When $\nabla_{1}\neq\nabla_{2}$, the larger gradient contributes more to the sum, so the computed weight is more reflective of the larger gradient. However, because $\omega_{p}^{c}$ is a scalar applied uniformly across the feature map, the distinct spatial contributions from $R_{1}$ and $R_{2}$ are merged. Consequently, the resulting attribution map $\Phi_{\mathrm{GradCAM}}^{c}$ loses the spatial granularity necessary to differentiate between the contributions of multiple objects.
II C Relevancy-propagation-based methods
II C 1 Background
Relevancy-propagation-based methods, such as LRP or DeepLIFT, explain the predictions of classifiers by back-propagating a relevancy score throughout the network, to attribute importance to each input feature. For relevancy-propagation-based explanation methods, we consider DeepLIFT [12] and LRP [11].
DeepLIFT [12] attributes the difference in the outputs $f_{c}(\bm{x})-f_{c}\left(\tilde{\bm{x}}\right)$ on the differences between the input image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ and a perturbation baseline (reference) $\tilde{\bm{x}}\;=\;\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$. Specifically, DeepLIFT runs forward propagation for both $\bm{x}$ and $\tilde{\bm{x}}$ inputs. Then, it stores the difference of the weighted activation $\Delta z_{s,t}=w_{s,t}^{(l,l+1)}\left(\bm{x}_{s}^{l}-\tilde{\bm{x}}^{l}_{s}\right)$, where $\bm{x}_{s}^{l}$ and $\tilde{\bm{x}}^{l}_{s}$ are the activations of the neuron $s$ in the $l$ layer and $w_{s,t}^{(l,l+1)}$ is the weight between neuron $s$ and neuron $t$ in the next layer $l+1$. Finally, the backpropagation of the difference in the outputs to the final input contributions $\Phi_{\text{DeepLift}}$ is achieved with the following rules:
$r_{t}^{(L)}=\begin{cases}f_{c}(\bm{x})-f_{c}(\tilde{\bm{x}}),&t=c,\\ 0,&t\neq c,\end{cases}$ (16)
and
$r_{s}^{(l)}=\sum_{t}\frac{\Delta z_{s,t}}{\displaystyle\sum_{s^{\prime}}\Delta z_{s^{\prime},t}}\;r_{t}^{(l+1)},$ (17)
and finally
$\Phi_{\text{DeepLift}}(f_{c},\bm{x})_{s}=r_{s}^{(1)}\,,$ (18)
where $L$ is the output layer and $r_{s}^{(1)}$ is the contribution of each feature on the input layer. For further propagation rules, we refer the reader to [12].
Layer Relevance Propagation (LRP)[11] assigns a relevance score $r$ to each input feature of a network by propagating the output of the model back through the network to the input layer. LRP interprets the neural network as a flow graph, redistributing prediction relevance backwards through the layers using specific propagation rules.
Relevance scores $r$ for a given layer $l$ are propagated to the neurons in the preceding layer as follows:
$r_{s}^{(l)}\;=\;\sum_{t}\frac{z_{s,t}}{\displaystyle\sum_{s^{\prime}}z_{s^{\prime},t}}\;r_{t}^{(l+1)}\,,$ (19)
where $z_{s,t}$ represents the contribution of neuron $s$ in layer $l$ to the activation of neuron $t$ in layer $l+1$, ensuring that total relevance is conserved across layers. LRP employs specific propagation rules to distribute relevance back to input features depending on the network architecture and task; see Montavon et al. [45] for details. For the final layer, the relevance $r_{t}^{(L)}$ is initialized as the output score for the target class $c$. The LRP method $\Phi_{\text{LRP}}$ for attributing the input space is defined as:
$\Phi_{\text{LRP}}\bigl{(}f_{c},\bm{x}\bigr{)}_{s}\;=\;r_{s}^{(1)}.$ (20)
Figure 3: LRP visualizations for an image from the DG-ML dataset. a) Original image showing two classes: Urban Land and agricultural Land; b) Pixel-wise reference map (blue: Urban Land; orange: agricultural Land); c) LRP explanation for agricultural Land (normalized range: 0–1, unnormalized range: 0–0.02); d) LRP explanation for Urban Land (normalized range: 0–1, unnormalized range: 0–0.18).
II C 2 Limitations
As DeepLIFT calculates relevance by comparing the output of a model for the input $\bm{x}$ and a reference input $\tilde{\bm{x}}$, expressed as $f_{c}(\bm{x})-f_{c}\left(\tilde{\bm{x}}\right)$, it encounters similar issues to perturbation-based methods (see Section  II A 2). Sturmfels et al. [46] investigate the impact of various reference baselines for the Integrated Gradients method, which also employs a difference-from-reference approach analogous to DeepLIFT. Their study visually examines reference baselines such as Gaussian blur, uniform values, blurring, maximum distance, and black baselines. They find that the selection of an appropriate baseline requires domain-specific knowledge to ensure it accurately represents missingness within the data domain. This challenge is amplified in RS image scene classification due to the availability of various spectral bands in the data.
A central property of LRP and related relevancy-propagation-based methods such as Integrated Gradients [47] is relevance conservation: the total relevance assigned to the input for class $c$ equals the network output, $\sum_{j}r_{j}^{c}=f_{c}(\bm{x})$ [45]. In natural images, discriminative information is typically concentrated within the object of interest. In contrast, RS scenes usually contain several LULC classes, so that a single class may occur in multiple, spatially disconnected regions; the discriminative information is therefore distributed rather than localized. Because the overall relevance budget is fixed, two classes that yield comparable output scores but occupy differently sized areas receive different relevance densities: the larger region is attributed a lower per-pixel relevance than the smaller region. Fig. 3 illustrates this issue using an image from the DG-ML dataset containing differently sized two classes: agricultural land ($c_{\text{major}}$) and urban land ($c_{\text{minor}}$). As the outputs of the model for both classes are similar $(f_{c_{\text{major}}}(\bm{x})\approxeq f_{c_{\text{minor}}}(\bm{x}))$ the maximum relevance to be propagated by LRP is similar. The higher number of relevant pixels for $c_{\text{major}}$ results in lower attributed relevance per pixel compared to $c_{\text{minor}}$. The $r^{c}_{j}$ for $c_{\text{major}}$ span values in the interval $[0,0.02]$, while the $r^{c}_{j}$ for $c_{\text{minor}}$ range within $[0,0.18]$. As attributions are typically normalized between 00 and $1$, this can produce attribution maps that are highly sensitive to outliers.
III Explanation metrics for RS image scene classification
In this section, we first provide a brief overview of the categorization of explanation metrics. We then introduce selected explanation metrics from different categories and discuss their suitability to estimate the quality of explanations for RS image scene classification. Explanation metrics offer a quantitative way to assess the quality of explanation methods. An explanation can be defined as a “presentation […] in human-understandable terms” [22]. Because this definition centers on subjective human interpretation, it inherently lacks a single, absolute ground truth. Therefore, it is not possible to estimate an explanation error, as one would estimate the performance error of a DL model [24]. Furthermore, explanation quality depends on more than the explanation method itself. When the decision-making process of a model is erroneous, even correctly derived explanations can appear inconsistent or counterintuitive, since explanations inherently reflect the internal logic of the model [48]. To address these challenges, researchers have proposed metrics across several categories that capture distinct aspects of explanation quality. In our study, we consider five different categories of explanation metrics: Faithfulness, Robustness, Localization, Complexity and Randomization. Formally, the quality of an explanation is assessed using a metric $\Psi$, which evaluates the effectiveness of an explanation method:
$\Psi(\Phi,\bm{x},f,\hat{\bm{y}})=\hat{q},$ (21)
where $\hat{q}$ is a scalar quality score. For simplicity, we omit explicit parameterization unless necessary.
III A Faithfulness
III A 1 Background
The first category that we consider is faithfulness, which measures how well explanations reflect the actual decision-making process of the model. A faithful explanation accurately reflects why the model made a specific prediction. Metrics in this category assess this alignment by evaluating whether features identified as important, significantly influence the prediction of the model. Let us consider $\bm{x}$ as the input image and $V\subseteq[H]\times[W]$ a set of pixels to remove. We define the perturbation function $\mathcal{P}_{\mathbb{X}}(\bm{x},V)$ which replaces each feature in $V$ with a predefined baseline value. The general faithfulness category $\Psi_{\mathrm{F}}(\bm{x},V)$ is defined as:
$\Psi_{\mathrm{F}}(\bm{x},V)=\bigl{\lvert}\,f(\bm{x})-f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr{)}\bigr{\rvert}.$ (22)
Perturbations are commonly applied to the most relevant pixels first (MoRF) or the least relevant ones first (LeRF).
To evaluate explanation methods for RS image scene classification, we consider two faithfulness metrics: i) Faithfulness Estimate (FE)[28], which measures the correlation between feature importance and prediction change, and ii) Iterative Removal Of Features (IROF)[29], which assesses prediction certainty decay when removing the most relevant segments.
Faithfulness Estimate (FE)[28] quantifies the alignment between attribution magnitude and model sensitivity by measuring the correlation between the total attribution on the top-$K$ pixels and the resulting drop in prediction when those pixels are removed. Formally, we define the FE metric $\Psi_{\mathrm{FE}}$ as:
$\Psi_{\mathrm{FE}}=\mathrm{corr}_{V:|V|=K}\Bigl{(}\sum_{(j,k)\in V}\Phi(\bm{x})_{j,k},\,f(\bm{x})-f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr{)}\Bigr{)}.$ (23)
Here, $\Phi(\bm{x})_{j,k}$ denotes the attributed relevancy at pixel $(j,k)$, and $V\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ is the set of the $K$ pixel indices with the largest attribution values. The function $\mathrm{corr}$ denotes the Pearson correlation coefficient. To compute $\Psi_{\mathrm{FE}}$, one first ranks all pixels by $\Phi(\bm{x})_{j,k}$, selects the top-$K$ indices $V$, and then correlates the summed attributions on $V$ with the change $f(\bm{x})-f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr{)}$. A higher value of $\Psi_{\mathrm{FE}}$ indicates that the explanation method attributes the behavior of the model, i.e., i.e., removing features deemed important by the method results in a significant drop in prediction certainty.
Iterative Removal Of Features (IROF)[29] measures how the prediction certainty changes when the most relevant features are removed but it considers the $K$ most relevant segments rather than pixels. Let us define $\{S_{m}\}_{m=1}^{M}$ as a partition of the image $\bm{x}$ into $M$ disjoint segments and sort them by their relevance, that:
$\frac{1}{|S_{1}|}\sum_{(j,k)\in S_{1}}\Phi(\bm{x})_{j,k}\;\geq\;\cdots\;\geq\;\frac{1}{|S_{M}|}\sum_{(j,k)\in S_{M}}\Phi(\bm{x})_{j,k}.$ (24)
For each $\kappa=0,1,\dots,K$, define the cumulative removal set $V_{\kappa}\;=\;\bigcup_{m=1}^{\kappa}S_{m},$ and the corresponding perturbed input $\mathcal{P}_{\mathbb{X}}(\bm{x},V_{\kappa}).$ Finally, the IROF score $\Psi_{\mathrm{IROF}}$ for $\bm{x}$ is obtained by normalizing each perturbed model output by the original confidence:
$\Psi_{\mathrm{IROF}}(\bm{x})=\frac{1}{K+1}\sum_{\kappa=0}^{K}\frac{f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V_{\kappa})\bigr{)}}{f(\bm{x})}\,.$ (25)
A higher $\Psi_{\mathrm{IROF}}$ indicates that removing highly attributed segments causes a rapid drop in prediction confidence, and thus reflects a more faithful explanation.
III A 2 Limitations
Faithfulness metrics inherit several limitations from perturbation-based explanation methods, as both rely on input perturbations to evaluate changes in the prediction certainty of a model. Consequently, they exhibit similar issues to those discussed in Section II A 2. For example, like perturbation-based methods, most faithfulness metrics require a carefully chosen baseline to avoid distribution shifts that can affect the predictions of the model.
In addition, the FE metric relies on a strong assumption: that removing a small number of high-attributed pixels will lead to a measurable change in model output. This assumption implicitly presumes that class-relevant information is concentrated in a small subset of pixels. However, RS scenes often contain multiple spatially disjoint regions associated with the same LULC class. These regions exhibit redundancy, as similar class-relevant features are repeated across the scene. In such cases, removing a small number of pixels does not produce a meaningful change in prediction, and the attribution sum remains nearly constant across different pixel subsets. As a result, the Pearson correlation used by FE may become numerically unstable, since both numerator and denominator can approach zero, leading to undefined or misleading scores even for faithful explanations.
In contrast to FE, some faithfulness metrics, such as IROF, evaluate the behavior of the model over a sequence of increasingly perturbed inputs. For these trajectory-based metrics, the order in which features are removed plays a critical role in shaping the evaluation outcome. To illustrate this, let us consider a scene in which a class $c_{\text{major}}$ occupies most of the image area, whereas a different class $c_{\text{minor}}$ covers only a small region. In such images, perturbing pixels belonging to $c_{\text{major}}$ do not significantly impact prediction certainty, as multiple pixels or regions contain similar predictive information. Consequently, metrics employing the MoRF strategy, which sequentially removes the most relevant pixels first, fail to capture meaningful changes in prediction certainty, as demonstrated by the slow decline of $f(\bm{x})$ in Fig. 4. In contrast, perturbations of the fewer, highly informative pixels associated with $c_{\text{minor}}$ lead to a rapid decrease in prediction certainty. This discrepancy results in biased metric outcomes, incorrectly suggesting that explanations for $c_{\text{minor}}$ are faithful, while explanations for $c_{\text{major}}$ appear less faithful. The LeRF strategy, which removes the least relevant pixels first, provides a better alternative to address this issue. By preserving the crucial predictive features of $c_{\text{minor}}$ until later perturbation steps, the LeRF strategy ensures a stable prediction certainty and yields a more balanced evaluation of explanation methods.
Figure 4: Comparison of MoRF and LeRF removal strategies for the target class $c$ (agricultural land, orange). Here, $K$ is the percentage of removed pixels and $f_{c}(\bm{x})$ represents the prediction certainty. Top row: a) Original image from the DG-ML dataset; b) Pixel-wise reference map. Agriculture land in orange; MoRF removal: c) $K=10\%$, $f_{c}(\bm{x})=0.98$; d) $K=50\%$, $f_{c}(\bm{x})=0.64$; e) $K=90\%$, $f_{c}(\bm{x})=0.31$. Bottom row: f) Grad-CAM attribution map for $c$; LeRF removal: g) $K=10\%$, $f_{c}(\bm{x})=0.99$; h) $K=50\%$, $f_{c}(\bm{x})=0.99$; i) $K=90\%$, $f_{c}(\bm{x})=0.72$.
III B Robustness
III B 1 Background
The second category we consider is robustness, which assesses the stability of explanations under small input perturbations. Similar, to adversarial attacks on DL models [49], explanations can also be fooled [50, 51]. Thus, ensuring that the explanation is invariant to minimal perturbations in the input space is crucial for critical applications. We require small perturbations $\delta$ such that $\left\|\delta\right\|_{2}<\varepsilon$ and expect $\left\|\mathcal{P}_{\mathbb{X}}(\bm{x})-\bm{x}\right\|_{2}<\varepsilon$ for a small positive constant $\varepsilon$, assuming that the model output approximately stayed the same $f(\bm{x})\approx f\left(\mathcal{P}_{\mathbb{X}}(\bm{x})\right)$. The category $\Psi_{\mathrm{ROB}}$ can be defined as:
$\Psi_{\mathrm{ROB}}(\Phi,\bm{x},\varepsilon)=\max_{\|\delta\|_{2}\leq\varepsilon}\bigl{\|}\,\Phi(\bm{x})\;-\;\Phi(\bm{x}+\delta)\bigr{\|}_{2}.$ (26)
To evaluate the robustness of explanation methods for RS image scene classification, we use i) Average Sensitivity (AS)[30], which measures sensitivity to small perturbations using a Monte Carlo-based approximation and ii) Local Lipschitz Estimate (LLE)[28], which estimates Lipschitz continuity to assess explanation stability under minor input changes.
Average Sensitivity (AS)[30] evaluates the sensitivity of the explanation method by perturbing samples using a Monte-Carlo-based approximation. The AS metric $\Psi_{\mathrm{AS}}$ is defined as:
$\Psi_{\mathrm{AS}}=\text{avg}_{\bm{x}+\delta\in\mathcal{N}_{\epsilon}(\bm{x})\leq\varepsilon}\left[\frac{\|\Phi(\bm{x})-\Phi(\bm{x}+\delta)\|}{\|\bm{x}\|}\right],$ (27)
where $\varepsilon$ defines the radius of a discrete, finite-sample neighborhood around each input sample $\bm{x}$. This neighborhood denoted as $\mathcal{N}_{\epsilon}(\bm{x})$, includes all samples that are within a distance of $\varepsilon$ from $\bm{x}$. A lower score is indicative of more robustness.
The Local Lipschitz Estimate (LLE)[28] estimates the Lipschitz continuity of the explanation, which measures how much the explanation changes with respect to the input under a minor perturbation, meaning the prediction did not change. The LLE metric $\Psi_{\mathrm{LLE}}$ is defined as:
$\Psi_{\mathrm{LLE}}=\max_{\bm{x}+\delta\in\mathcal{N}_{\epsilon}(\bm{x})\leq\epsilon}\left[\frac{|\Phi(\bm{x})-\Phi(\bm{x}+\delta)|_{2}}{|\bm{x}-(\bm{x}+\delta)|_{2}}\right],$ (28)
where lower values indicate less change concerning the change in input, which is desirable.
III B 2 Limitations
A key hyper-parameter in this metric category is the noise type. If the introduced noise does not naturally occur in the acquired images, it may generate OoD samples, reducing the reliability of the model and thereby affecting the validity of the explanations. When evaluating the robustness of explanation methods, it is important to introduce noise that does not result in OoD samples. This consideration also depends on the model architecture, as many models exhibit inherent resilience to certain types of noise. For instance, in SAR images, multiplicative noise such as speckle is pervasive [52]. In contrast, for multispectral images, noise is typically independent of the data and generally additive in nature, such as additive white Gaussian noise [53, 54].
III C Localization
III C 1 Background
The third category we consider is localization, which measures the spatial accuracy of explanations by comparing them to reference maps or bounding boxes. This comparison can be quantified using a similarity measure SIM between the generated explanation map $\Phi(\bm{x})$ and a pixel-level reference map $\bm{s}_{gt}$. The category $\Psi_{\mathrm{LOC}}$ is defined as:
$\Psi_{\mathrm{LOC}}(\Phi,\bm{x},\bm{s}_{gt},\text{SIM})=\text{SIM}\left(\Phi(\bm{x}),\bm{s}_{gt}\right),$ (29)
where SIM can be a measure such as Intersection over Union, Pearson correlation, or SSIM depending on the nature of the task and reference map.
To evaluate the localization capabilities of explanation methods for RS image scene classification, we use i) Top-K Intersection (TKI)[32], which measures the fraction of top-ranked attributions within the pixel-level reference map, and ii) Relevance Rank Accuracy (RRA)[31], which assesses the concentration of high-relevance values within the reference map.
Top-K Intersection (TKI), as proposed by Theiner et al. [32], measures the fraction of the $K$ highest ranked attributions that lie on the binary reference mask $\bm{s}_{gt}$. Let us define
$I_{K}=\{\,i_{1},i_{2},\ldots,i_{K}\}\quad\text{and}\quad S_{gt}=\{\,p\mid\bm{s}_{gt}(p)=1\},$ (30)
where $I_{K}$ is the set of the top-$K$ ranked pixel indices and $S_{gt}$ the set of true-positive pixels in the mask. The intersection of these sets is:
$\bm{s}_{1-K}=I_{K}\;\cap\;S_{gt}.$ (31)
The TKI metric $\Psi_{\mathrm{TKI}}$ is then defined as:
$\Psi_{\mathrm{TKI}}=\frac{\bigl{|}\bm{s}_{1-K}\bigr{|}}{K}.$ (32)
This measures how many of the $K$ most important features, according to $\bm{e}$, match the pixel‐level reference set $S_{gt}$. If all top-$K$ pixels are within the reference mask, then $\Psi_{\mathrm{TKI}}=1$, therefore a higher score is better.
The RRA [31] measures the concentration of high-intensity relevance values within the pixel-level reference map. Let $|S_{\text{tot}}|$ be the size of the pixel-level reference map mask. Select the $K$ highest relevance values, according to $\Phi(\bm{x})$, count how many of these values lie within the actual pixel locations, and divide by $|S_{\text{tot}}|$. This can be expressed as:
$P_{\text{top }K}=\left\{p_{1},p_{2},\ldots,p_{K}\mid R_{p_{1}}>R_{p_{2}}>\ldots>R_{p_{K}}\right\},$ (33)
where $P_{\text{top }K}$ is the set of pixels with relevance values $R_{p_{1}},R_{p_{2}},\ldots,R_{p_{K}}$ sorted in decreasing order until the $K$-th pixel. The RRA metric $\Psi_{\mathrm{RRA}}$ is then computed as:
$\Psi_{\mathrm{RRA}}=\frac{\left|P_{\text{top }K}\cap S_{\text{tot}}\right|}{|S_{\text{tot}}|}.$ (34)
III C 2 Limitations
In RS scene classification, label dependencies arise when the presence or absence of one label influences the likelihood of another, which is a common scenario in RS where certain objects (e.g., ships) typically appear in specific environments (e.g., sea). These dependencies are not considered by localization metrics. Furthermore, the reliability of localization metrics heavily depends on accurate and up-to-date pixel-level reference maps. Obtaining such annotations is costly and prone to errors. Inaccuracies or obsolescence in the reference maps can compromise the localization reliability, potentially leading to incorrect assessments. Additionally, most localization metrics become unreliable, when the target class spans a large area of the input image. Let us consider the TKI metric in the scenario where we have a single label for the input image. Here, the binary reference mask $\mathbf{s}_{gt}$ for the target label becomes a matrix of ones, encompassing all pixels. Thus, $\Psi_{\mathrm{TK}}$ inherently achieves its maximum value of 1, regardless of the explanation $\bm{e}$. Since $\mathbf{s}_{gt}$ includes every pixel, the intersection $\mathbf{s}_{1-K}=\mathbf{r}_{1-K}\cap\mathbf{s}_{gt}$ reduces to $\mathbf{r}_{1-K}$ itself, as all top-ranked indices in $\mathbf{r}_{1-K}$ necessarily lie within $\mathbf{s}_{gt}$. Consequently, the numerator and denominator of $\Psi_{\mathrm{TK}}$ both equal $K$, yielding $\Psi_{\mathrm{TK}}=1$. In this case, the metric cannot differentiate meaningful explanations from arbitrary ones. This phenomenon is well-documented for simpler localization metrics [55], such as the Pointing Game [56]. However, in RS scene classification this also creates challenges for more sophisticated metrics, as in most RS scene classification tasks, as classes tend to span larger areas.
III D Complexity
III D 1 Background
The fourth category we consider is complexity, which quantifies the simplicity and interpretability of explanations, ensuring that they are useful for human understanding. These metrics evaluate properties such as sparsity, entropy, or other measures of complexity in the explanation map $\Phi(\bm{x})$. The category $\Psi_{\mathrm{CMP}}$ is defined as:
$\Psi_{\mathrm{CMP}}(\Phi,f,\bm{x},C)=C\left(\Phi(\bm{x})\right),$ (35)
where $C$ represents a complexity measure, chosen based on the context of interpretability. To evaluate the complexity of explanations for RS image scene classification, we use i) Sparseness (SP)[33] and ii) Complexity (CO)[34].
Sparseness (SP)[33] is a method for evaluating the sparsity of explanations and is defined as the Gini index of the explanation. It is calculated by summing the product of the ranks of the input features and their attributions and dividing by the sum of the attributions. The SP metric $\Psi_{\mathrm{SP}}$ is defined as:
$\Psi_{\mathrm{SP}}=\frac{\displaystyle\sum_{j=1}^{H}\sum_{k=1}^{W}\bigl{(}2\,\rho_{j,k}-D-1\bigr{)}\,\Phi(\bm{x})_{j,k}}{D(D-1)\,\displaystyle\sum_{j=1}^{H}\sum_{k=1}^{W}\Phi(\bm{x})_{j,k}},$ (36)
where $\hat{\bm{e}}_{j,k}$ is the attribution of pixel $\bm{x}_{j,k}$, $\rho_{j,k}$ is the rank of $\Phi(\bm{x})_{j,k}$ among all $D$ attributed input features when sorted in non‐decreasing order, $H$ and $W$ denote the image height and width, i.e. $D=H\times W$. A higher $\Psi_{\mathrm{SP}}$ indicates a sparser explanation.
Complexity (CO)[34] is defined using the Shannon entropy calculation, which measures the amount of uncertainty or randomness in the attribution map. It is calculated by summing the product of the probabilities of the attributions and the logarithm of the probabilities of the attributions. The CO metric $\Psi_{\mathrm{CO}}$ is defined as:
$\Psi_{\mathrm{CO}}=-\sum_{j=1}^{H}\sum_{k=1}^{W}p_{j,k}\,\ln\bigl{(}p_{j,k}\bigr{)},$ (37)
where
$p_{j,k}=\frac{|\Phi(\bm{x})_{j,k}|}{\sum_{a=1}^{H}\sum_{b=1}^{W}|\Phi(\bm{x})_{a,b}|}$ (38)
is the fractional contribution of attribution $\Phi(\bm{x})_{j,k}$ , and a higher $\Psi_{\mathrm{CO}}$ indicates greater uncertainty (complexity). A uniformly distributed attribution would have the highest possible complexity score.
Figure 5: Explanations for a DG-ML image labeled as Forest and their corresponding SP metric scores. a) Original image; b) Pixel-wise reference map (green: Forest); c) Occlusion explanation ($\Psi_{\mathrm{SP}}=0.08$); d) LRP explanation ($\Psi_{\mathrm{SP}}=0.47$).
III D 2 Limitations
Explanation metrics in the complexity category generally assume that an explanation is simpler if it attributes relevance to fewer features for a given prediction. This assumption originates from explaining predictions for natural images, where a class is often defined by a few unique features. There, only a small fraction of the input features contribute to the final prediction. However, classification of RS images often relies more on spectral and texture features, a larger set of repetitive features is relevant, which is overlooked by complexity metrics.
A key limitation of the Gini-based SP metric $\Psi_{\mathrm{SP}}$ is its tendency to penalize explanations with uniformly distributed attributions. In a scenario where a class spans nearly the entire image, a comprehensive explanation assigns relevance to multiple regions exhibiting discriminative spectral or textural signals. This uniform distribution is interpreted by $\Psi_{\mathrm{SP}}$ as high complexity, thereby misclassifying an accurate explanation as less concise. In contrast, an explanation for a minor class confined to a single region is favored with a low complexity score, even though such concentration does not necessarily improve interpretability. This limitation is illustrated in Fig. 5, which shows two explanations for a scene labeled as forest, a class that spans nearly the entire image. The attribution map from the Occlusion method (Fig. 5(c)) uniformly highlights all forest pixels, yielding a low SP value ($\Psi_{\mathrm{SP}}=0.08$). Although most of the input pixels are attributed as relevant, the map remains easy to interpret. By contrast, the LRP map (Fig. 5(d)) concentrates relevance on a small, localized area, producing a higher sparsity score ($\Psi_{\mathrm{SP}}=0.47$). Here the SP metric favors a compact but misleading attribution, mistaking spatial concentration for interpretive simplicity.
III E Randomization
III E 1 Background
The fifth category we consider is randomization, which tests the reliability of explanations by introducing randomness into model parameters or prediction targets. These metrics evaluate the consistency of explanations under such perturbations to ensure that explanations are not artifacts of specific parameter settings or configurations. The category $\Psi_{\mathrm{RAND}}$ is defined as:
$\Psi_{\mathrm{RAND}}(\Phi,f,\bm{x},\mathcal{R},\text{SIM})=\text{SIM}\left(\Phi(\bm{x}),\Phi(\bm{x},f,\mathcal{R})\right),$ (39)
where $\mathcal{R}$ represents a randomization operation (e.g., model parameter shuffling or target randomization) and SIM measures the similarity between the original explanation $\Phi(\bm{x},f)$ and the explanation generated under the randomized condition $\Phi(\bm{x},f,\mathcal{R})$. To evaluate the randomization robustness of explanation methods for RS image scene classification, we use i) Efficient Model Parameter Randomization Test (MPRT)[57], which measures explanation complexity under model parameter randomization and ii) Random Logit (RL)[36], which evaluates stability by comparing explanations for target and non-target classes using Structural Similarity Index Measure (SSIM).
Efficient Model Parameter Randomization Test (MPRT)[57] measures the relative rise in the complexity of the explanation $\Phi\bigl{(}\bm{x},f_{\mathcal{R}}\bigr{)}$ from a fully randomized model $f_{\mathcal{R}}$ for image $\bm{x}$. The MPRT metric $\Psi_{MPRT}$ is defined as:
$\Psi_{MPRT}=\frac{\xi\bigl{(}\Phi\bigl{(}\bm{x},f_{\mathcal{R}}\bigr{)}\bigr{)}-\xi\bigl{(}\Phi\bigl{(}\bm{x},f\bigr{)}\bigr{)}}{\xi\bigl{(}\Phi\bigl{(}\bm{x},f\bigr{)}\bigr{)}}.$ (40)
Here, $\xi:\mathbb{R}^{D}\!\to\!\mathbb{R}$ denotes a complexity function. To define $\xi$, a histogram‐entropy measure grounded in Shannon entropy is employed. Specifically, the vectorized attributions $\bm{e}=\{\,a_{j,k}\}_{j,k}$ are placed into $B$ distinct bins; for the $d$-th bin, the frequency $\eta_{d}$ is computed and the corresponding normalized probability $\pi_{d}$ is
$\xi(\bm{e})=-\sum_{d=1}^{B}\pi_{d}\,\log\!\bigl{(}\pi_{d}\bigr{)},\quad\text{where}\quad\pi_{d}=\frac{\eta_{d}}{\sum_{d^{\prime}=1}^{B}\eta_{d^{\prime}}}\,.$ (41)
The Random Logit (RL)metric [36] is defined as the SSIM over the explanation of the reference label and an explanation of non-target class $\bm{y}^{\prime}$. The RS metric $\Psi_{\mathrm{RL}}$ is defined as:
$\Psi_{\mathrm{RL}}=\operatorname{SSIM}\left(\Phi(\bm{x},f,\hat{\bm{y}}),\Phi\left(\bm{x},f,\bm{y}^{\prime}\right)\right),$ (42)
where $\Phi(\bm{x},f,\hat{\bm{y}})$ is the explanation generated for the prediction $\hat{\bm{y}}$ and $\Phi\left(\bm{x},f,\bm{y}^{\prime}\right)$ is the explanation generated for a non-target class $\bm{y}^{\prime}$. Lower values indicate that the explanations are not correlated, which is desirable.
III E 2 Limitations
Due to their minimal sensitivity to variations in input data characteristics, randomization metrics are also robust for RS image classification. Metrics like MPRT focus solely on how the randomization of model weights affects the outcomes of the explanation method, isolating the influence of the parameters the model has learned. In contrast, RL measures the similarity between explanations for a target class and a non-target class. However, this metric may encounter challenges with label dependencies. For example, if the target class and the non-target class are both influenced by a shared present class, regions associated with the shared class may exhibit high relevance for both classes. In such cases, the attribution similarity can be disproportionately high for both target and non-target classes, because the shared class dominates the explanation outcomes.
IV Dataset Description and Experimental Design
IV A Dataset Description
The experiments are conducted using three datasets: 1) BigEarthNet-S2 v2.0 (BEN-v2.0)[58], 2) DeepGlobe Multi-Label (DG-ML)[59, 60], and 3) Five Billion Pixel Multi-Label (FBP-ML)[61]. Table I summarizes the key characteristics of the considered datasets.
TABLE I: Summary of key attributes of utilized RS datasets, including the number of images ($|D|$), unique labels ($L$), average labels per image, number of bands ($C$), image sizes and the type of pixel-level reference maps.
Dataset $|\mathcal{D}|$     $L$ Avg. $L$     $C$ Image Size (Spatial Resolution) Pixel-Level Reference Maps
per Image
BEN-v2.0 $250.249$ $19$ $2.95$ $10$ 120$\times$120 (10 m) Thematic Product
DG-ML $18.185$ $6$ $1.71$ $3$ 120$\times$120 (0.5 m) Manually Annotated
FBP-ML $109.200$ $24$ $2.61$ $4$ 256$\times$256 (4 m) Manually Annotated
BEN-v2.0 [58] is a multi-label dataset which extends the BigEarthNet archive [62] and covers ten countries in Europe. We use the Sentinel-2 imagery with 12 bands and 10 m spatial resolution. Each $1200$ m×$\times 1200$ m tile is subdivided into $120\times 120$-pixel patches. Class annotations are taken from the CORINE Land-Cover Map inventory 2018, assigning subsets of 19 LULC classes to 549 488 patches. We adopt the official train–validation–test split defined in [58].
DG-ML is a multi-label dataset, which Burgert et al. [60] derived from the original DeepGlobe Land-Cover Classification Challenge [59]. The original DeepGlobe corpus comprises 1,949 manually annotated RGB tiles of size $2448\times 2448$ pixels at 0.5 m spatial resolution, collected over Thailand, Indonesia, and India. The LULC classes are urban land, agricultural land, range land, forest, water, barren land, and unknown. Each tile was divided into a grid of $120\times 120$-pixel patches, and multi-labels were assigned from the corresponding pixel-level reference maps, while patches containing the class “unknown” were discarded. Only 20 % of single-class patches and all multi-class patches are retained, yielding 30 443 samples split into training (60%), validation (20%) and test (20%) subsets.
FBP-ML is a multi-label dataset derived from the Five Billion Pixels (FBP) collection [61], comprising 150 Gaofen-2 RGB–NIR scenes (4 m spatial resolution) that together cover roughly 50 000 km2 in China. Following the procedure used for DG-ML, each scene is partitioned into a grid of $256\times 256$-pixel patches and labeled with a 24-class scheme that distinguishes artificial, agricultural, and natural surfaces. The final dataset consists of 87 360 training samples (80 %), 10 920 validation samples (10 %), and 10 920 test samples (10 %).
IV B Evaluation Protocol
In the experiments, we evaluate explanation methods and metrics for feature attribution in RS image scene classification using the pipeline illustrated in Fig. 1. We generate explanations by applying different feature attribution methods to the prediction of DL models for the considered RS datasets. Next, we estimate the quality of these explanations using the considered explanation metrics. Then, the reliability of these metrics is assessed by using the MetaQuantus [24] framework.
The MetaQuantus framework [24] considers that the quality of an explanation cannot be objectively measured due to the lack of ground truth. It tests the explanation metrics against two failure modes: Noise Resilience (NR)of a metric, which is the resilience against minor perturbations, and Adversary Reactivity (AR), which is the reactivity of the metrics to disruptive perturbations. A perturbation is considered minor when the predicted class label does not change and disruptive when the predicted class label changes. To test against these failure modes, two consistency criteria Intra-Consistency (IAC)and Inter-Consistency (IEC)are utilized.
The IAC criterion measures whether the explanation scores remain similarly distributed after minor perturbations and whether they become dissimilarly distributed following disruptive perturbations. Let $\hat{\mathbf{q}}\in\mathbb{R}^{N}$ be the unperturbed metric scores for $N$ samples and $\mathbf{q}^{\prime}_{k}\in\mathbb{R}^{N}$ the corresponding perturbed scores for $k=1,\dots,K$. The IAC score is defined as:
$\text{IAC}\;=\;\frac{1}{K}\,\sum_{k=1}^{K}\,d\bigl{(}\hat{\mathbf{q}},\mathbf{q}^{\prime}_{k}\bigr{)},$ (43)
where $d$ is the Wilcoxon signed-rank test and $0\leq\text{IAC}\leq 1$. A higher value indicates that unperturbed and perturbed metric scores are more similarly distributed.
The IEC criterion evaluates the consistency of metrics in preserving ranking order across multiple explanation methods. Let $\bar{\mathbf{Q}}\in\mathbb{R}^{N\times L}$ be the unperturbed metric scores for $N$ samples and $L$ explanation methods, averaged over $K$ perturbations. Additionally, let $\overline{\mathbf{Q}}^{\prime}\in\mathbb{R}^{N\times L}$ be the corresponding averaged perturbed metric scores. We define a binary ranking agreement matrix $\mathbf{U}^{t}\in[0,1]^{N\times L}$, where $t\in\{M,D\}$ indicates a minor ($M$) or disruptive ($D$) perturbation. The IEC score is defined as:
$\text{IEC}=\frac{1}{N\times L}\sum_{i=1}^{N}\sum_{j=1}^{L}U_{i,j}^{t},$ (44)
with $0\leq\text{IEC}\leq 1$. The entries $U_{i,j}^{t}$ capture how consistently the ranking of the metrics agrees between unperturbed $\bar{\mathbf{Q}}$ and perturbed $\overline{\mathbf{Q}}^{\prime}$. For minor perturbations ($t=M$) the ordering of the $L$ methods for each sample must remain identical, whereas for disruptive perturbations ($t=D$) the unperturbed score must always ranked higher than its perturbed counterpart.
The Meta-Consistency (MC)score combines both intra- and inter-consistency criteria across different failure modes as:
$\mathbf{MC}\;=\;\left(\frac{1}{|\bm{m}^{*}|}\right)\bm{m}^{*T}\bm{m},$ (45)
where
$\bm{m}\;=\;\begin{bmatrix}\mathrm{IAC}_{NR}\\ \mathrm{IAC}_{AR}\\ \mathrm{IEC}_{NR}\\ \mathrm{IEC}_{AR}\end{bmatrix}\quad\text{and}\quad\bm{m}^{*}\;=\;\mathbb{1}^{4},$ (46)
with the all-ones vector representing an ideal explanation metric that achieves a value of 1 on all criteria. A metric with an ${MC}$ score close to 1 indicates strong performance across all criteria tested with higher values reflecting greater overall robustness and reliability. The MC score is averaged over two types of perturbation: model perturbation and input perturbation. In model perturbation, random noise is added to the parameters, whereas in input perturbation, it is added to the input data.
IV C Experimental Setup
We have systematically evaluated ten explanation metrics across five widely used feature attribution methods on multi-label RS datasets (BEN-v2.0, DG-ML, and FBP-ML). To this end, we trained a ResNet-50 model on BEN-v2.0 and ResNet-18 models on DG-ML and FBP-ML, subsequently computing explanations for the respective model predictions. The ResNet architecture was selected due to its robust performance and demonstrated efficacy when paired with popular explanation methods.
We evaluated the quality of the explanation using the considered metrics and evaluated their reliability with MetaQuantus [24] on 512 randomly drawn samples, a sample size demonstrated to be sufficient in previous work [24]. The hyperparameter were set as in the original publication ($k=5$, three iterations). Gaussian noise was chosen for the input perturbation to preserve physical consistency, in line with the recommendations in [63]. The choice of model depth does not substantially affect the outcome of the evaluation, as the performance of MetaQuantus has been shown to be largely invariant with respect to model architecture and depth [24]. For model training, we used the AdamW optimizer, initialized at a learning rate of $10^{-2}$ and a weight decay of $0.9$. All experiments were performed on Nvidia H100 80GB GPUs.
We limit our analysis to five explanation methods, as prior work [24] indicates a minimal influence of the number of methods on the result scores. Additionally, we used a Random baseline explanation method that samples values randomly from a uniform distribution $U(0,1)$. For a multi-label case with a label set $\{1,…,K\}$, we apply the same method separately to each label $c\in\{1,…,K\}$ to obtain a feature attribution map for each class $c$. Furthermore, as common in the literature, we summed the feature attributions over the channels.
The hyperparameters for the explanation methods were configured as follows. For DeepLIFT, the perturbation baseline value was set to 00. For GradCAM, the last convolutional layer was analyzed and the explanations were upscaled by bilinear interpolation. The LIME method employed a Euclidean exponential kernel (kernel width 500) as a similarity measure, used linear regression as an interpretable model, and generated 15 superpixels using SLIC. For LRP, following [45], the first third of the network was initialized with the LRP-$\gamma$ rule, the second third with the LRP-$\epsilon$ rule and the final third with the LRP-00 rule. Lastly, the Occlusion method employed a sliding window of size $(50,50)$ with strides of $(10,10)$ across the channels and the perturbation baseline was set to 00. The quality of the explanations was assessed using 10 explanation metrics implemented via the Quantus framework [25] with default hyperparameters. Quantus is recognized as the most suitable framework for evaluating explanations [64]. The explanation metrics were computed over 1024 randomly drawn test samples from each dataset. The explanation methods are implemented using the Captum framework [65] with modifications to Captum, Quantus, and MetaQuantus to support multi-label image scene classification.
V Experimental Results
We carried out different kinds of experiments in order to i) assess the reliability of explanation metrics for RS image scene classification problems, and ii) evaluate the effectiveness of feature attribution methods based on the most reliable explanation metrics. This ordering ensures that the selection of suitable explanation methods is grounded in metrics whose reliability has been rigorously validated.
V A Evaluation of the considered explanation metrics
Figure 6: Mean MC scores for the considered explanation metrics: a) scores associated to all RS datasets (DG-ML, FBP-ML and BEN-v2.0); b) scores associated to each individual dataset (DG-ML, FBP-ML, BEN-v2.0). Metrics are categorized into five color-coded groups: Complexity (brown), Faithfulness (blue), Localization (orange), Randomization (green), and Robustness (red).
In Fig. 6(a), we report the mean MC score of each explanation metric, where the scores have been averaged over all three multi‑label RS datasets (BEN-v2.0, DG-ML, and FBP-ML), resulting in a single bar per metric. In contrast, Fig. 6(b) decomposes these results by dataset: for each explanation metric, three separate bars show its mean MC score on DG-ML, FBP-ML and BEN-v2.0 individually. Across both figures, the magnitude of the metrics remains stable, indicating that metric reliability is robust whether computed on the combined dataset or on each dataset in isolation.
From Fig. 6(a) one can see that the metrics from the robustness category are the most reliable for RS image scene classification. In particular, LLE achieves the highest mean MC score of $0.79$, while AS attains the second-highest score of $0.68$. This is because the metrics in this category have fewer limitations compared to other categories and are therefore particularly well-suited for RS image scene classification (see Section III B 2). By observing the figure, one can also see that the localization metrics, RRA and TKI, perform worst with mean MC scores of $0.52$ and $0.51$, respectively. The primary cause of the low performance of localization metrics arises from their methodological constraint: They are unreliable when there is a single LULC class present in $\bm{x}$.
On a closer look, the evaluation of the input perturbation test on these metrics (TKI in Fig. 7(a) and RRA in Fig. 7(b)) reveals consistent performance patterns. Specifically, the input perturbation test yields similar trends across the four evaluation scores that contribute to the MC score (see Eq. 46). Both TKI and RRA exhibit high values for $\text{IAC}_{\text{NR}}$ and relatively high $\text{IEC}_{\text{NR}}$ values. On the other hand, both metrics show low values for $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$.
Figure 7: Results of the evaluation of the input test on the DG-ML dataset for the localization metrics: a) Top-K Intersection and b) Relevance Rank Accuracy.
The NR scores indicate consistency over minor perturbations, while the AR scores reflect changes in response to disruptive perturbations. The pattern observed in these scores responds to the methodological flaws of the localization metrics described in Section III C 2. Specifically, when a single LULC class spans the entire image, the explanation coincides exactly with the reference map. Therefore, the metric always yields an output value of $1$. Since the $\text{IAC}_{\text{ER}}$ and $\text{IEC}_{\text{ER}}$ scores measure the metrics consistency over minor perturbations, both scores will consistently assign the highest score of 1. However, as the output is constant, disruptive perturbations cannot change the output of the metric, leading to low $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$ values. This flaw is particularly visible in the DG-ML dataset, as it has the lowest average amount of LULC classes per image of the considered datasets.
Furthermore, it is visible in Fig. 6(a) that the randomization metrics also perform comparatively well with the second highest average scores. This is in accordance with the assumption that the metrics of this category perform independently of the data. From the figure, one can see that the MPRT metric achieves a mean MC score of $0.68$, while the RL metric attains a score of $0.60$.
Figure 8: Results of the evaluation of the input test on the DG-ML dataset for the complexity metrics: a) Complexity; and b) Sparseness.
The mean MC values for the metrics CO and SP in the complexity category are $0.63$ and $0.57$, as visualized in Fig. 6(a). Similar to the localization metrics, the complexity metrics also struggle with the characteristic of RS image data in the case that single LULC classes can span large parts of an image, yet only a small region provides the discriminative information necessary for accurate classification. In this case, the metrics tend to attribute higher complexity to the explanation. This deficit is again reflected in the low $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$ values, as shown for the CO metric in Fig. 8(a) and the SP metric in Fig. 8(b).
Figure 9: Results of the evaluation of the input test on the BEN-v2.0 dataset for the faithfulness metrics: a) Faithfulness Estimate; and b) Iterative Removal Of Features.
As shown in Fig. 6(a), the mean MC scores for the faithfulness category are $0.56$ and $0.60$ for FE and IROF, respectively. Fig. 9(a) and Fig. 9(b) present the detailed evaluation results for the BEN-v2.0 dataset, which we consider as the perturbation baseline is expected to have a stronger impact for multi-spectral RS images. Similar to the localization metrics, the $\text{IEC}_{\text{AR}}$ score is notably low, with values of 0.28 for FE and 0.01 for IROF. This aligns with methodological findings that faithfulness metrics face similar challenges as localization metrics when an image contains redundant image features for classes that span large parts of the input image.
In such cases, the prediction certainty is not expected to decrease significantly when important features are removed (see Section III A 2). A low $\text{IEC}_{\text{AR}}$ indicates that the metric scores for the perturbed input are consistently lower than for the unperturbed input. However, when the prediction of the model remains unchanged, faithfulness metrics still assign a high score, keeping the score on the unperturbed input high and consequently leading to low $\text{IEC}_{\text{AR}}$ values. Conversely, the $\text{IAC}_{\text{AR}}$ scores for FE and IROF are high, with values of $0.82$ and $1.0$, respectively. This suggests that the distribution of scores remains unchanged even under disruptive perturbations. One possible explanation is a distribution shift induced by OoD samples that occurs due to the perturbation, which prevents the faithfulness scores from effectively capturing the expected changes.
Figure 10: Scores of the IROF metric on the BEN-v2.0 dataset under four perturbation configurations: (i) mean-based LeRF, (ii) mean-based MoRF, (iii) black-based MoRF, and (iv) uniform-based MoRF.
To further investigate how perturbation baselines and removal strategies influence the reliability of faithfulness metrics on multispectral images, we conduct an ablation study using the IROF metric with four configurations: mean-based LeRF, mean-based MoRF, black-based MoRF, and uniform-based MoRF (Fig.10). The results show that combining the mean baseline with LeRF yields the highest reliability (0.75), followed by mean-based MoRF (0.64), black-based MoRF (0.61), and uniform-based MoRF (0.58). As discussed in Section III A 2, removing the most relevant pixels first can fail to capture meaningful changes in prediction certainty, particularly when large image regions contain similar predictive information. In contrast, the LeRF approach preserves critical features until later perturbation steps, leading to a more balanced evaluation. Moreover, employing a mean-value baseline reduces the likelihood of generating OoD samples, further enhancing the reliability of the faithfulness scores.
V B Evaluation of the considered feature attribution methods
TABLE II: Average metric scores obtained by the considered explanation methods for BEN-v2.0
Faithfulness $\uparrow$ Robustness $\uparrow$ Localization $\uparrow$ Complexity $\uparrow$ Randomization $\uparrow$
Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
DeepLIFT 0.66 0.59 0.49 1.00 0.49 0.48 0.60 0.30 0.40 0.49
GradCAM 0.73 0.56 0.88 0.99 0.50 0.43 0.46 0.47 0.81 0.67
LIME 0.61 0.54 0.50 0.05 0.49 0.44 0.58 0.46 0.66 0.26
LRP 0.61 0.51 0.71 0.99 0.48 0.45 0.55 0.20 0.23 0.25
Occlusion 0.64 0.48 0.87 0.71 0.49 0.44 0.35 0.52 0.74 0.42
Random baseline 0.39 0.44 0.85 0.60 0.47 0.42 0.07 0.00 0.96 0.26
TABLE III: Average metric scores obtained by the considered explanation methods for DG-ML
Faithfulness $\uparrow$ Robustness $\uparrow$ Localization $\uparrow$ Complexity $\uparrow$ Randomization $\uparrow$
Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
DeepLIFT 0.67 0.65 0.41 1.00 0.62 0.45 0.71 0.48 0.46 0.66
GradCAM 0.65 0.65 0.68 0.96 0.74 0.48 0.43 0.42 0.83 0.49
LIME 0.53 0.66 0.66 0.56 0.64 0.44 0.52 0.50 0.72 0.23
LRP 0.69 0.66 0.61 0.99 0.64 0.47 0.54 0.26 0.22 0.42
Occlusion 0.75 0.68 0.69 0.43 0.69 0.47 0.31 0.40 0.82 0.34
Random baseline 0.44 0.57 0.84 0.78 0.59 0.37 0.19 0.03 0.97 0.36
TABLE IV: Average metric scores obtained by the considered explanation methods for FBP-ML
Faithfulness $\uparrow$ Robustness $\uparrow$ Localization $\uparrow$ Complexity $\uparrow$ Randomization $\uparrow$
Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
DeepLIFT 0.75 0.59 0.45 1.00 0.46 0.49 0.70 0.43 0.29 0.69
GradCAM 0.75 0.62 0.75 0.99 0.48 0.50 0.57 0.85 0.71 0.46
LIME 0.65 0.55 0.56 0.41 0.45 0.45 0.55 0.51 0.67 0.16
LRP 0.69 0.58 0.53 0.99 0.43 0.47 0.65 0.28 0.29 0.26
Occlusion 0.81 0.58 0.69 0.81 0.46 0.48 0.50 0.48 0.73 0.25
Random baseline 0.37 0.39 0.85 0.45 0.45 0.41 0.06 0.00 0.98 0.17
In this subsection, we assess the considered feature attribution methods for RS image scene classification using the considered explanation metrics grouped by category. Tables II, III, and IV report the average metric scores from each dataset. Faithfulness is assessed using the FE and IROF metrics. As observed in the tables, Occlusion achieves the highest FE score for the DG-ML and FBP-ML datasets, whereas GradCAM performs best for the BEN-v2.0 dataset. For the IROF metric, DeepLIFT, Occlusion, and GradCAM attain the highest scores for BEN-v2.0, DG-ML and FBP-ML, respectively. Averaged across all datasets, Occlusion outperforms all other methods in the FE metric with an average score of 0.73, while GradCAM and DeepLIFT achieve the highest IROF score of 0.61. Both metrics exhibit low variance across datasets. Robustness is measured using the AS and LLE metrics. In the AS metric, the random baseline outperforms all methods on the DG-ML and FBP-ML datasets. This is because the AS metric assesses how much the explanation changes on average when noise is introduced. Given a sufficiently large sample size, a uniformly drawn distribution remains relatively stable under such perturbations, leading to higher scores for the random baseline. Since the LLE metric measures the maximum change, the random baseline performs worse. Among explanation methods, GradCAM achieves the highest AS score for BEN-v2.0 and FBP-ML, while Occlusion performs best for DG-ML. For the LLE metric, DeepLIFT achieves the highest score across all datasets, closely followed by GradCAM and LRP. Both perturbation-based methods (LIME and Occlusion) perform worse in the LLE metric.
Localization is evaluated using the TKI and RRA metrics. Across all datasets and metrics, except for RRA on BEN-v2.0, GradCAM performs best. The variance across the scores for the different datasets and methods remains low, except for the TKI metric in the DG-ML dataset, which exhibits higher variability. Complexity is assessed using the SP and CO metrics. For the SP metric, DeepLIFT achieves the highest performance across all datasets. The results for the CO metric vary across datasets; averaged over all datasets, GradCAM yields the best performance with an average score of 0.58. The random baseline performs poorly, as expected, given that the CO metric is defined as the entropy of the explanation, which reaches its maximum when the explanation is uniformly distributed. Randomization is quantified using the RL and MPRT metrics. The random baseline achieves the highest RL score across all datasets, with an average performance of 0.97, aligning with the findings of [66], which demonstrate that randomized attribution maps can achieve high scores in top-down randomization metrics. While this outcome exposes a limitation of the RL metric, it remains valuable in identifying explanation methods that are largely independent of the model and therefore unreliable. For the MPRT metric, GradCAM achieves the best result on the BEN-v2.0 dataset, while DeepLIFT performs best on DG-ML and FBP-ML.
As illustrated in Fig. 11, the largest dataset-dependent differences occur in the TKI metric, which quantifies how well explanations align with pixel-level reference maps. Specifically, the scores are notably higher for the DG-ML dataset, likely due to its low number of unique labels and lower average number of labels per image, compared to the other datasets (see Table I).
Fig. 12(a) displays the overall average performance of the evaluated explanation methods across all considered metrics. The results show that GradCAM achieves the highest score ($0.65$), followed by DeepLIFT ($0.58$) and Occlusion ($0.57$). LRP and LIME achieve slightly lower scores of $0.52$ and $0.50$, respectively, while the Random Baseline performs worst at $0.46$. In contrast, Fig. 12(b) presents the weighted average scores of the most reliable metrics. Here, GradCAM again performs best ($0.66$), but DeepLIFT achieves a higher relative score ($0.63$), surpassing Occlusion ($0.52$). Notably, LRP improves in ranking, reaching a score of $0.53$, while LIME and the Random Baseline exhibit a more pronounced decline, with scores of $0.43$ and $0.37$, respectively. Comparing both figures, GradCAM consistently outperforms other methods, maintaining the highest rank across both evaluation approaches. However, the usage of more reliable metrics in Fig. 12(b) results in an increased performance gap between methods, with DeepLIFT improving its relative standing and LIME and the Random Baseline exhibiting a sharper decline. This shows that using more reliable metrics leads to a more accurate evaluation and confirms the robustness of GradCAM. The larger reduction in Random Baseline performance indicates that the most reliable metrics for RS image scene classification effectively penalize random explanations.
Figure 11: Comparison of metric averages across datasets. Each metric is color-coded by category.
Figure 12: Comparison of explanation methods based on evaluation metrics. (a): normalized average of all metrics across all RS datasets; (b): normalized average of the best-performing metric (CO, IROF, LLE, RRA, MPRT) for each category across all RS datasets.
Overall, GradCAM emerges as the most versatile explanation method, as it achieves strong performance across multiple evaluation categories. It excels in faithfulness, localization, and randomization, making it particularly well-suited for RS image scene classification.
VI Conclusion and Guidelines
In this study, for the first time in RS, we have systematically investigated the suitability of explanation methods and metrics for RS image scene classification by combining methodological analysis with extensive experiments on three datasets. In particular, we analyzed five widely used feature attribution methods: Occlusion [9], LIME [10], LRP [11], DeepLIFT [12] and GradCAM [13] and ten explanation metrics spanning five categories: faithfulness (FE [28], IROF [29]), robustness (AS [30], LLE [28]), localization (RRA [31], TKI [32]), complexity (SP [33], CO [34]) and randomization-based (MPRT [35], RL [36]).
Our methodological analysis shows that perturbation‐based approaches, such as Occlusion and LIME, are critically dependent on the choice of perturbation baseline. In RS scenes, replacing spectral content with a fixed value can generate OoD samples. Furthermore, the redundancy of features across large regions hinders the perturbation function to significantly decrease the prediction of a model, which lowers the reliability of perturbation-based methods. Gradient‐propagation methods like GradCAM are limited in their ability to distinguish between multiple instances of the same object in a scene. Relevancy-propagating methods like LRP tend to underestimate the contribution of LULC classes that span larger areas of a scene. Regarding evaluation metrics, our findings indicate that faithfulness metrics inherit the limitations of perturbation‐based methods and need to be parametrized accordingly. In contrast, robustness metrics demonstrate greater stability. Localization metrics require accurate pixel-level reference maps. Furthermore, localization and complexity metrics are unreliable when a single class spans a large part of the scene. Randomization metrics are often independent of the data. To support our methodological findings, we conducted two sets of experiments on three RS datasets: BEN-v2.0, DG-ML, and FBP-ML. In the first set of experiments, we meta-evaluated ten explanation metrics to assess their reliability. We observe that metrics from the robustness and randomization categories performed well. Conversely, as expected from the methodological analysis, metrics from the faithfulness and localization categories are more unreliable. In the second set of experiments, we use the most reliable explanation metrics to determine which explanation methods are best suited for RS image scene classification. Among the evaluated methods, GradCAM emerged as the most robust feature attribution method, demonstrating relatively stable performance across different datasets and evaluation metrics. However, it is important to note that no single method performed optimally across all evaluation criteria.
Following our methodological and experimental analysis, we develop guidelines for selecting suitable explanation metrics and explanation methods for scene classification in RS:
  1. 1.
    We recommend using robustness and randomization metrics due to their reliability across datasets and explanation methods.
  2. 2.
    Faithfulness metrics should be used cautiously, as they depend on perturbation baselines and may not reliably reflect model behavior. To mitigate this issue, we recommend using the LeRF removal strategy, which helps to maintain reliability by gradually removing less relevant features first. Furthermore, we recommend utilizing the mean value as a perturbation baseline.
  3. 3.
    We suggest using localization metrics only when accurate pixel-level reference maps are available. In particular, we recommend using more advanced localization metrics that also account for the size of the target class. Complexity metrics should be interpreted carefully, as they depend strongly on the spatial extent of the target class. Similarly to localization metrics, complexity metrics could also be modified to account for the target class size.
  4. 4.
    We recommend using GradCAM as a general baseline due to its versatility and ease of interpretation. Among the evaluated methods, it consistently achieves high scores across multiple metric categories.
As a final remark, we would like to note that, due to the lack of ground truth on explanations, it remains not possible to fully validate an explanation metric [24]. The quantifiable reliability of a metric does not necessarily indicate its validity [66]. Thus, each evaluation must be critically examined. Future work should focus on refining explanation metrics specifically designed for RS image scene classification. In particular, developing perturbation strategies that better accommodate spectral variability and feature redundancy may enhance the reliability of faithfulness metrics and, by extension, perturbation-based explanation methods. Although our study concentrated on feature attribution methods commonly used in RS (while being originally designed for natural images), a promising direction for future research is to develop and evaluate explanation techniques tailored specifically to the characteristics of RS images.

### v3

@@
 In recent years, deep learning (DL) has emerged as the dominant approach for scene classification problems due to the high capability of DL models to extract and exploit the complex spatial and spectral content of RS images.
-RS image scenes can contain multiple land-use land-cover (LULC) classes and thus can be simultaneously associated with different class labels (i.e.
-multi-labels).
+RS image scenes can contain multiple land-use land-cover (LULC) classes and thus can be simultaneously associated with different class labels (i.e., multi-labels).
 Accordingly, the existing models are in general capable of assigning multiple LULC class labels to each RS image scene in an archive [1, 2].
 Despite their success, DL models often remain enigmatic black boxes, with almost no human understanding of their internal processes [3].
-This lack of transparency raises concerns as decision-makers increasingly rely on these black box systems [4].
+This lack of transparency raises concerns as decision-makers increasingly rely on these black-box systems [4].
 New regulations for critical applications, such as the General Data Protection Regulation [5] and the Artificial Intelligence Act [6], require interpretability to ensure transparent decisions.
 Beyond regulatory concerns, effective interpretability can also improve model performance by addressing issues such as limited or biased training data, outliers, vulnerability to adversarial attacks, and overfitting [4].
-To this end, explainable artificial intelligence (xAI)has emerged as a research area focused on revealing the factors behind DL model predictions.
+To this end, explainable artificial intelligence (xAI) has emerged as a research area focused on revealing the factors behind the predictions of the DL model.
 Although xAI methods are widely used in the CV community [7], they have only recently gained attention in RS.
-For a comprehensive overview of xAI in RS we refer the reader to [8].
+For a comprehensive overview of xAI in RS, we refer the reader to [8].
 The most prominent type of explanation methods for RS scene classification problems are feature attribution methods [8].
@@
 [8] reviewed approximately 200 RS publications, finding that the vast majority evaluated qualitative, while only 16 studies included quantitative evaluation using explanation metrics.
-This reliance on qualitative assessment raises concerns, as studies from the CV community have shown that qualitative assessment is insufficient for robust verification of explanation methods [21, 22].
-One of the key reasons for the limited use of quantitative evaluations in RS is the absence of clear guidelines for assessing explanation quality in this domain.
-The need for such guidelines is underscored by the inherent difficulty of quantitative evaluation since explanation metrics depend heavily on their parametrization [23, 22].
+This reliance on qualitative assessment raises concerns, as studies in the CV community have shown that qualitative assessment is insufficient for the robust verification of explanation methods [21, 22].
+One of the key reasons for the limited use of quantitative evaluations in RS is the absence of clear guidelines to assess the quality of the explanation in this domain.
+The need for such guidelines is underscored by the inherent difficulty of quantitative evaluation, since the explanation metrics depend heavily on their parameterization [23, 22].
 Moreover, the lack of ground truth for the explanations renders them unverifiable [24], requiring the use of a wide range of metrics for a more comprehensive evaluation [25].
-In many RS studies only a small part of this range is analyzed, often focusing on single aspects, such as localization capability [17], while only a few studies have evaluated a broader variety of metrics [26, 27].
+In many RS studies, only a small part of this range is analyzed, often focusing on single aspects, such as localization capability [17], while only a few studies have evaluated a broader variety of metrics [26, 27].
 In particular, none of these studies have addressed the fact that most explanation metrics have been originally designed for natural images and can be limited when applied to RS scene classification.
@@
 These methods and metrics are selected for their popularity in RS and to cover a broad spectrum of evaluation criteria.
-We conduct a methodological analysis to examine whether existing explanation methods and metrics are suitable for RS image scene classification.
-Then, we perform an extensive experimental evaluation using MetaQuantus [24] to assess the reliability and effectiveness of explanation metrics under minor perturbations and disruptive changes, supporting our methodological findings with experimentally validation.
+We perform a methodological analysis to examine whether the existing explanation methods and metrics are suitable for the classification of RS images.
+Then, we perform an extensive experimental evaluation using MetaQuantus [24] to assess the reliability and effectiveness of the explanation metrics under minor perturbations and disruptive changes, supporting our methodological findings with experimental validation.
 We would like to note that, unlike previous work that primarily surveys the state of the art in xAI for RS [8], our study goes beyond a methodological overview by conducting a systematic, in-depth analysis of explanation methods and evaluation metrics.
@@
   • Based on our methodological and experimental findings, we develop practical guidelines to assist RS practitioners in selecting appropriate explanation methods and evaluation metrics for RS image scene classification and recommend suitable explanation methods and metrics.
-Our pipeline can be reproduced and adapted for different datasets or classification tasks in RS, providing a valuable resource for the community.
+Our pipeline can be reproduced and adapted for different datasets or classification tasks in RS, providing a valuable resource for the RS community.
 The remainder of this paper is organized as follows.
 Section II and Section III provide a methodological evaluation of the effectiveness of explanation methods and explanation metrics for RS image scene classification, respectively.
-In Section IV we describe the datasets and experimental setup.
+In Section IV we describe the datasets and the experimental setup.
 The results of our experiments are presented in Section V.
@@
 Figure 1: The evaluation protocol for selecting appropriate explanation methods and metrics for RS image scene classification.
-It involves generating explanations using various feature attribution methods, estimating their quality using selected metrics, assessing the reliability of the metric with MetaQuantus, and using the most reliable metrics to select the most suitable explanation methods.
+It involves generating explanations using various feature attribution methods, estimating their quality using selected metrics, assessing the reliability of the metric with MetaQuantus [24], and using the most reliable metrics to select the most suitable explanation methods.
 II Explanation methods for RS image scene classification
-In this section, we methodologically analyze the explanation methods to address the unique challenges posed by the RS images.
+In this section, we provide a methodological analysis of the considered explanation methods with respect to the specific challenges of RS images.
 Most xAI research has focused on natural images, which differ significantly from RS images [19].
-To this end, we first analyze a range of explanation methods commonly applied in RS scene classification regarding these differences.
-We focus on local post-hoc explanation methods that generate feature attribution maps, as they are most commonly used in RS [8].
-Feature attribution methods can be further grouped into backpropagation-based and perturbation-based methods [37].
-Perturbation-based methods alter model inputs to observe changes in the model output, while backpropagation methods propagate values from the model output to the input space, typically gradients or relevancy scores.
-We consider three categories: i) perturbation-based methods, ii) relevancy-propagation-based methods and iii) gradient-propagation-based methods.
-To provide a comprehensive overview, our analysis includes five popular feature attribution methods covering all categories: 1) two perturbation-based methods (Occlusion [9], LIME [10]); 2) two relevancy-propagation-based methods (LRP [11], DeepLIFT [12]) and 3) one gradient-propagation-based method (GradCAM [13]).
+In this paper, we examine commonly applied explanation methods in RS scene classification, focusing on methods that generate feature attribution maps, since these methods are most widely used in RS [8].
+However, we note that alternative paradigms such as Multiple Instance Learning [37, 38, 39] or counterfactual explanations [40] have also been utilized in RS.
+Feature attribution methods can be grouped into perturbation-based and backpropagation-based methods [41].
+Perturbation methods modify inputs to measure output changes, while backpropagation methods propagate values from outputs to inputs, typically gradients or relevancy scores.
+We consider three categories: i) perturbation-based, ii) relevancy-propagation-based, and iii) gradient-propagation-based methods.
+Our analysis covers five representative methods: 1) two perturbation-based methods (Occlusion [9], LIME [10]); 2) two relevancy-propagation-based methods (LRP [11], DeepLIFT; [12]) and 3) one gradient-propagation-based method (GradCAM [13]).
 For each category, we first provide a general description, followed by a definition of the explanation methods and a discussion of the limitations considering the differences between RS and natural images.
-Let us consider a RS image scene classification problem where a DL model $f$ is trained on a RS image archive $\bm{X}_{\mathrm{tr}}$ made up of $N$ pairs: $(\bm{x}_{i},y_{i})_{i=1}^{N}$ for single‐label classification, or $(\bm{x}_{i},\bm{y}_{i})_{i=1}^{N}$ for multi‐label classification.
-For simplicity, we omit the single-label notation.
-Each image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ represents a scene captured by an airborne or satellite sensor, where $C$ denotes the spectral bands and $H,W$ correspond to spatial dimensions.
-The model $f$ maps an input image $\bm{x}$ to a predicted set of labels:
-$f(\bm{x})=\hat{\bm{y}}$ (1)
-where $\hat{\bm{y}}$ represents the predicted class probabilities or binary assignments.
+Let us consider a RS scene classification problem, where a DL model $f$ is trained on an archive $\bm{X}_{\mathrm{tr}}$ of $N$ pairs $(\bm{x}_{i},\bm{y}_{i})_{i=1}^{N}$ for multi-label classification (the single-label case is omitted for simplicity).
+Each image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ represents a scene captured by an airborne or satellite sensor, with $C$ spectral bands and $H,W$ spatial dimensions.
+The model $f$ maps an input image to predicted labels $f(\bm{x})=\hat{\bm{y}}$, where $\hat{\bm{y}}$ represents the predicted class probabilities or binary assignments.
 To interpret the reasoning behind the prediction of a model $f$ for an input $\bm{x}$, post-hoc explanation methods, such as feature attribution methods $\Phi$, can be applied:
@@
 For simplicity, we omit explicit parameterization unless necessary.
-II A Perturbation-based methods
-II A 1 Background
-Perturbation-based explanation methods explain model predictions by measuring changes in the output between the original and perturbed inputs.
-Let $\mathbb{X}\subseteq\mathbb{R}^{H\times W\times C}$ be the space of valid inputs (e.g., images) and let $\Theta$ be a set of perturbation parameters (e.g., masks, noise levels, patch locations).
-A perturbation function $\mathcal{P}_{\mathbb{X}}$ is defined as:
-$\mathcal{P}_{\mathbb{X}}:\mathbb{X}\times\Theta\;\to\;\mathbb{X},$ (3)
-such that for any $\bm{x}\in\mathbb{X}$ and any $\theta\in\Theta$, the output
-$\tilde{\bm{x}}\;=\;\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$ (4)
-remains in $\mathbb{X}$.
+II.A Perturbation-based methods
+II.A.1 Background
+Perturbation-based methods explain predictions by measuring output changes between original and perturbed inputs.
+Let $\mathbb{X}\subseteq\mathbb{R}^{H\times W\times C}$ denote the input space and $\Theta$ a set of perturbation parameters.
+A perturbation function $\mathcal{P}_{\mathbb{X}}:\mathbb{X}\times\Theta\;\to\;\mathbb{X}$ is defined as:
+$\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)\;=\;\tilde{\bm{x}}$ (2)
+with $\bm{x}\in\mathbb{X}$ and $\theta\in\Theta$.
 In other words, $\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$ modifies $\bm{x}$ according to the perturbation parameter $\theta$, while still preserving the overall structure such that $\tilde{\bm{x}}$ is a valid point in the same space.
-We define the category of perturbation-based methods $\Phi_{\text{Perturbation}}$ as:
-$\Phi_{\text{Perturbation}}(f,\bm{x},\mathcal{P}_{\mathbb{X}},\theta)=|f(\bm{x})-f(\mathcal{P}_{\mathbb{X}}(\bm{x},\theta))|.$ (5)
-Common perturbation functions are blurring [38], introducing noise [30], or replacing parts of the image [39]. For perturbation-based explanation methods, we consider Occlusion [9] and LIME [10].
-The Occlusion method [9] is defined as follows.
-Let each $\theta\in\Theta$ specify a region (mask) $\Omega\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ and a constant perturbation baseline $\alpha\in\mathbb{R}$.
-We define the perturbation function:
+Common perturbation functions are blurring [42], introducing noise [30], or replacing parts of the image [43].
+We consider Occlusion [9] and LIME [10] as perturbation-based methods.
+In the Occlusion [9] method, each $\theta\in\Theta$ specifies a region (mask) $\Omega\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ and a baseline $\alpha\in\mathbb{R}$. The perturbation function is:
 $\bigl[\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)\bigr]_{j,k}\;=\;\begin{cases}\alpha,&\text{if }(j,k)\in\Omega,\\ \bm{x}_{j,k},&\text{otherwise}.\end{cases}$ (3)
@@
 $\Phi_{\text{Occlusion}}(j,k)\;=\;\frac{1}{\bigl|\{\theta:(j,k)\in\Omega\}\bigr|}\sum_{\theta:(j,k)\in\Omega}\Delta s_{c}(\theta).$ (5)
-The LIME $\Phi_{\mathrm{LIME}}$ method explains a black-box model $f$ by locally approximating it with a simple, interpretable model trained on binary-encoded perturbations of the original input.
+The LIME method $\Phi_{\mathrm{LIME}}$ explains a black-box model $f$ by locally approximating it with an interpretable model trained on binary-encoded perturbations of the input.
 Let $S=\{S_{1},S_{2},\dots,S_{M}\}$ be a segmentation of $\bm{x}$ into $M$ superpixels.
-Denote by $\bm{z}\in\{0,1\}^{M}$ a binary vector that indicates which superpixels are present (1) or not present (0).
-We define the perturbation function:
+We define $\bm{z}\in\{0,1\}^{M}$ as a binary vector indicating the presence (1) or absence (0) of superpixels.
+The perturbation function is:
 $\mathcal{P}_{\mathbb{X}}:\mathbb{X}\times\{0,1\}^{M}\;\to\;\mathbb{X},$ (6)
@@
 $\mathcal{P}_{\mathbb{X}}(\bm{x},\bm{z})_{j,k}\;=\;\begin{cases}[\bm{x}]_{j,k},&\text{if }(j,k)\in S_{m}\text{ ,}\bm{z}_{m}=1,\\ \alpha,&\text{if }(j,k)\in S_{m}\text{ , }\bm{z}_{m}=0,\end{cases}$ (7)
-and $\alpha\in\mathbb{R}^{C}$ is a constant perturbation baseline.
-By toggling each superpixel via $\bm{z}_{m}\in\{0,1\}$, we create perturbed samples in the neighborhood of $\bm{x}$.
-LIME estimates a simple, interpretable model $g$ (e.g., a sparse linear model) that approximates $f$ on these perturbations near $\bm{x}$.
-Formally, let:
-$\Phi_{\mathrm{LIME}}(\bm{x})\;=\;\operatorname{argmin}_{g\in\mathcal{G}}\Bigl{[}\mathcal{L}\bigl{(}f,g,\Pi_{x}\bigr{)}\;+\;\Upsilon(g)\Bigr{]}.$ (11)
-Here, $\Upsilon(g)$ is a regularization term, such as the number of nonzero weights in a linear model or the depth of a decision tree, $\mathcal{L}$ is a loss function, and $\Pi_{x}(\tilde{\bm{x}})$ denotes a proximity measure that quantifies the distance from a perturbation $\tilde{\bm{x}}$ to $\bm{x}$.
-II A 2 Limitations
-A key component of perturbation-based methods is the perturbation baseline, which is intended to represent the absence of a feature.
-However, defining an appropriate perturbation baseline is non-trivial.
-Many RS studies simply employ a scalar value $\alpha$, often set to zero, such as in the quantitative evaluation of Kakageorgiou et al.
+where $\alpha\in\mathbb{R}^{C}$ is a constant baseline. Varying $\bm{z}_{m}$ generates perturbed samples around $\bm{x}$. Then, we fit a simple model $g$ that approximates $f$ near $\bm{x}$:
+$\Phi_{\mathrm{LIME}}(\bm{x})\;=\;\operatorname{argmin}_{g\in\mathcal{G}}\Bigl[\mathcal{L}\bigl(f,g,\Pi_{x}\bigr)\;+\;\Upsilon(g)\Bigr],$ (8)
+[…]
+II.A.2 Limitations
+A key component of perturbation-based methods is the perturbation baseline, which represents the absence of the occluded input feature.
+In practice, many RS studies use a scalar value $\alpha$, often zero, as in Kakageorgiou et al.
 [26].
-However, perturbations with a scalar perturbation baseline can yield Out-of-Distribution (OoD)samples [40].
-This is problematic because DL models are prone to unreliable predictions when the input does not reside in their learned distribution [41].
-An example case is shown in Fig.
-2(a), which depicts an RGB image from the DeepGlobe Multi-Label (DG-ML)dataset.
+However, scalar baselines can produce Out-of-Distribution (OoD) samples [44], and DL models yield unreliable predictions on inputs outside their training distribution [45].
+Fig.
+2(a) illustrates this limitation with an RGB image from the DeepGlobe Multi-Label (DG-ML) dataset [46].
 A rectangular patch is perturbed with the Occlusion method, using a black perturbation baseline ($\alpha=0$).
 Despite occluding a part of the class, the prediction of the model for the class agricultural land increased from $f_{c}(\bm{x})=0.90$ on the original input to $f_{c}(\tilde{\bm{x}})=0.99$ after the occlusion.
-Even when most of the pixels from the original image are perturbed, as shown in Fig.
-2(b), the model predicts with high certainty $f_{c}(\tilde{\bm{x}})=0.38$, highlighting the unpredictable behavior introduced by a poorly chosen perturbation baseline.
-This challenge is amplified in multispectral RS images, where different spectral bands (e.g., near-infrared or thermal) capture distinct information.
-Common baselines, such as all-zero vectors, can be unsuitable because certain bands rarely contain zero values in operational conditions.
-For example, in thermal imagery, a zero-valued perturbation baseline is not meaningful, as it does not reflect any typical temperature reading.
-Similarly, for a near-infrared band, some level of infrared radiation is almost always present, making an all-zero replacement an inadequate representation of the absence of a feature.
-Consequently, selecting a meaningful perturbation baseline requires more sophisticated approaches that account for the different spectral bands.
-Figure 2: Visualizations of perturbed images using $\alpha=0$ from the DG-ML dataset for the target class $c$ (agricultural land) and its prediction certainty for the original image: $f_{c}(\bm{{x}})=0.9$.
+Even with most pixels perturbed (Fig.
+2(b)), the model outputs $f_{c}(\tilde{\bm{x}})=0.38$, highlighting the unpredictable behavior introduced by a poorly chosen perturbation baseline.
+This challenge is amplified in multispectral RS images, where spectral bands capture distinct information.
+Common baselines, such as all-zero vectors, are unsuitable because some bands rarely take zero values in operational conditions.
+For example, in thermal images, a zero baseline has no physical meaning, and in near-infrared bands, some radiation is always present.
+Meaningful baselines are required to account for spectral characteristics.
+Figure 2: Visualizations of perturbed images using $\alpha=0$ from the DG-ML [46] dataset for the target class $c$ (agricultural land) and its prediction certainty for the original image: $f_{c}(\bm{{x}})=0.9$.
 a) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.99$: b) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.38$.
-Furthermore, the regions within each RS image are associated with different LULC classes, and unlike in natural images, background content does not exist.
-Let us consider the case where we have a LULC class $c$ present in the image and let $A_{c}(\bm{x})$ represent the total number of pixels in $\bm{x}$ corresponding to class $c$ (which can be interpreted as the discrete spatial extent of class $c$ within the image).
-If the area covered by the perturbation mask $A(\Omega)$ of the Occlusion method is smaller than $A_{c}(\bm{x})$, the perturbation may fail to obscure a sufficient number of relevant pixels to significantly decrease the prediction certainty.
-This is depicted in Fig.
-2(a), where we see a rectangular region that is occluded, which does not reduce $f_{c}(\bm{\tilde{x}})$.
-This limitation is less pronounced in natural images, where discriminative features are often unique (e.g., silhouette, object parts).
-In such cases, occluding typically leads to a noticeable reduction in prediction certainty even if $A_{c}(\bm{x})>A(\Omega)$.
+[…]
 A similar problem occurs for the LIME method, where $\mathcal{P}_{\mathbb{X}}$ partitions $\bm{x}$ in different superpixels and partially occludes them.
-Here, the prediction remains unchanged, unless most of the superpixels are replaced with the perturbation baseline (see Fig 2(b)), resulting in a linear model trained primarily on positive class predictions.
+Unless most superpixels are replaced (Fig.
+2(b)), the prediction remains unchanged, resulting in a linear model trained primarily on positive class predictions.
 This imbalance causes the estimated weights of the linear model to be unreliable, leading to an inaccurate interpretation of feature importance.
-Since no single superpixel significantly alters the decision of the model, LIME assigns almost equal importance to all superpixels.
-However, this relevance is lower than it should be, as the method underestimates the contribution of discriminative features due to the regularization term, potentially leading to wrong interpretations of the reasoning of the model.
+Since no single superpixel strongly influences the model prediction, LIME assigns nearly equal importance across superpixels, underestimating discriminative features due to regularization.
 Additionally, the more pixels are replaced by the perturbation baseline, the higher the proximity measure $\Pi_{x}$ due to the large distance from the original image.
 As a result, individual pixels are incorrectly attributed as less relevant.
-II B Gradient-propagation-based methods
-II B 1 Background
-Gradient-propagation-based methods utilize gradient information to generate visual explanations for deep neural network predictions.
-The gradient-propagation-based explanation method we examine is GradCAM [13], which incorporates gradient information flowing into any convolutional layer, typically the last one due to its high-level semantics and spatial information retention.
-The attribution map for a specific class $c$ is computed by first capturing the gradients of the class score $Y^{c}$ with respect to the activations $A^{p}$ of convolutional layer $l$.
-The importance weights $\omega^{c}_{p}$ for each feature map are determined by globally averaging the gradients over all spatial locations $(j,k)$:
-$\omega^{c}_{p}=\frac{1}{Z}\sum_{j,k}\frac{\partial Y^{c}}{\partial A^{p}_{j,k}}.$ (12)
-These weights represent a partial linearization of the network and denote the importance of each feature map $p$ for the target class $c$.
-A weighted combination of these maps followed by a ReLU ensures that only positive influences on the class prediction are visualized:
+II.B Gradient-propagation-based methods
+II.B.1 Background
+Gradient-propagation-based methods use gradients to explain model predictions.
+We consider GradCAM [13], which utilizes gradients flowing into a convolutional layer, typically the last one, due to its semantic and spatial content.
+For class $c$, the gradients of the class score $Y^{c}$ with respect to activations $A^{p}$ of layer $l$ yield importance weights $\omega^{c}_{p}$:
+$\omega^{c}_{p}=\frac{1}{Z}\sum_{j,k}\frac{\partial Y^{c}}{\partial A^{p}_{j,k}},$ (9)
+where global averaging collapses spatial locations $(j,k)$.
+These weights approximate the contribution of each feature map $p$ to class $c$.
+The attribution map is the weighted feature maps followed by ReLU, ensuring only positive contributions are visualized:
 $\Phi_{\mathrm{GradCAM}}^{c}=\text{ReLU}\left(\sum_{p}\omega^{c}_{p}A^{p}\right).$ (10)
-II B 2 Limitations
-Gradient-propagation-based methods exhibit limitations when applied to RS imagery, particularly when multiple LULC classes occur in a scene.
-For instance, in a RS image containing five airplanes, GradCAM localizes only a single instance [42], underscoring its inability to resolve multiple objects of interest at varying scales [43, 44].
+II.B.2 Limitations
+Gradient-propagation-based methods exhibit limitations when applied to RS scenes with multiple LULC classes.
+For instance, in a scene containing five airplanes, GradCAM localizes only a single instance [47], underscoring its inability to resolve multiple objects of interest at varying scales [48, 49].
 This failure originates from the reliance of GradCAM on a global gradient averaging mechanism, which collapses spatial gradient information into a single channel-wise importance weight.
@@
 When $\nabla_{1}\neq\nabla_{2}$, the larger gradient contributes more to the sum, so the computed weight is more reflective of the larger gradient.
-However, because $\omega_{p}^{c}$ is a scalar applied uniformly across the feature map, the distinct spatial contributions from $R_{1}$ and $R_{2}$ are merged.
-Consequently, the resulting attribution map $\Phi_{\mathrm{GradCAM}}^{c}$ loses the spatial granularity necessary to differentiate between the contributions of multiple objects.
-II C Relevancy-propagation-based methods
-II C 1 Background
-Relevancy-propagation-based methods, such as LRP or DeepLIFT, explain the predictions of classifiers by back-propagating a relevancy score throughout the network, to attribute importance to each input feature.
-For relevancy-propagation-based explanation methods, we consider DeepLIFT [12] and LRP [11].
-DeepLIFT [12] attributes the difference in the outputs $f_{c}(\bm{x})-f_{c}\left(\tilde{\bm{x}}\right)$ on the differences between the input image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ and a perturbation baseline (reference) $\tilde{\bm{x}}\;=\;\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$.
-Specifically, DeepLIFT runs forward propagation for both $\bm{x}$ and $\tilde{\bm{x}}$ inputs.
-Then, it stores the difference of the weighted activation $\Delta z_{s,t}=w_{s,t}^{(l,l+1)}\left(\bm{x}_{s}^{l}-\tilde{\bm{x}}^{l}_{s}\right)$, where $\bm{x}_{s}^{l}$ and $\tilde{\bm{x}}^{l}_{s}$ are the activations of the neuron $s$ in the $l$ layer and $w_{s,t}^{(l,l+1)}$ is the weight between neuron $s$ and neuron $t$ in the next layer $l+1$.
-Finally, the backpropagation of the difference in the outputs to the final input contributions $\Phi_{\text{DeepLift}}$ is achieved with the following rules:
+Since $\omega_{p}^{c}$ is a scalar applied uniformly to $A^{p}$, spatially distinct contributions from $R_{1}$ and $R_{2}$ merge, and the attribution map $\Phi_{\mathrm{GradCAM}}^{c}$ loses the granularity needed to separate multiple objects.
+II.C Relevancy-propagation-based methods
+II.C.1 Background
+Relevancy-propagation-based methods explain predictions by back-propagating relevance scores to attribute importance to input features. We consider DeepLIFT [12] and LRP [11].
+[…]
 $r_{t}^{(L)}=\begin{cases}f_{c}(\bm{x})-f_{c}(\tilde{\bm{x}}),&t=c,\\ 0,&t\neq c,\end{cases}$ (13)
@@
 $r_{s}^{(l)}=\sum_{t}\frac{\Delta z_{s,t}}{\displaystyle\sum_{s^{\prime}}\Delta z_{s^{\prime},t}}\;r_{t}^{(l+1)},$ (14)
-and finally
+and the final attribution $\Phi_{\text{DeepLift}}$ is defined as:
 $\Phi_{\text{DeepLift}}(f_{c},\bm{x})_{s}=r_{s}^{(1)},$ (15)
 where $L$ is the output layer and $r_{s}^{(1)}$ is the contribution of each feature on the input layer. For further propagation rules, we refer the reader to [12].
-Layer Relevance Propagation (LRP)[11] assigns a relevance score $r$ to each input feature of a network by propagating the output of the model back through the network to the input layer.
-LRP interprets the neural network as a flow graph, redistributing prediction relevance backwards through the layers using specific propagation rules.
-Relevance scores $r$ for a given layer $l$ are propagated to the neurons in the preceding layer as follows:
+LRP [11] propagates output relevance back to inputs by interpreting the network as a flow graph. For a layer $l$, relevance scores $r$ are redistributed as follows:
 $r_{s}^{(l)}\;=\;\sum_{t}\frac{z_{s,t}}{\displaystyle\sum_{s^{\prime}}z_{s^{\prime},t}}\;r_{t}^{(l+1)}\,$, (16)
-where $z_{s,t}$ represents the contribution of neuron $s$ in layer $l$ to the activation of neuron $t$ in layer $l+1$, ensuring that total relevance is conserved across layers.
-LRP employs specific propagation rules to distribute relevance back to input features depending on the network architecture and task; see Montavon et al.
-[45] for details.
-For the final layer, the relevance $r_{t}^{(L)}$ is initialized as the output score for the target class $c$.
-The LRP method $\Phi_{\text{LRP}}$ for attributing the input space is defined as:
+where $z_{s,t}$ is the contribution of neuron $s$ in layer $l$ to neuron $t$ in $l+1$, ensuring conservation of total relevance.
+Specific propagation rules depend on the model architecture and task [50].
+For the final layer, $r_{t}^{(L)}$ is initialized as the target class score $f_{c}(\bm{x})$.
+The input attribution LRP method $\Phi_{\text{LRP}}$ is then:
 $\Phi_{\text{LRP}}\bigl(f_{c},\bm{x}\bigr)_{s}\;=\;r_{s}^{(1)}.$ (17)
-Figure 3: LRP visualizations for an image from the DG-ML dataset.
-a) Original image showing two classes: Urban Land and agricultural Land; b) Pixel-wise reference map (blue: Urban Land; orange: agricultural Land); c) LRP explanation for agricultural Land (normalized range: 0–1, unnormalized range: 0–0.02); d) LRP explanation for Urban Land (normalized range: 0–1, unnormalized range: 0–0.18).
-II C 2 Limitations
-As DeepLIFT calculates relevance by comparing the output of a model for the input $\bm{x}$ and a reference input $\tilde{\bm{x}}$, expressed as $f_{c}(\bm{x})-f_{c}\left(\tilde{\bm{x}}\right)$, it encounters similar issues to perturbation-based methods (see Section  II A 2).
-Sturmfels et al.
-[46] investigate the impact of various reference baselines for the Integrated Gradients method, which also employs a difference-from-reference approach analogous to DeepLIFT.
-Their study visually examines reference baselines such as Gaussian blur, uniform values, blurring, maximum distance, and black baselines.
-They find that the selection of an appropriate baseline requires domain-specific knowledge to ensure it accurately represents missingness within the data domain.
-This challenge is amplified in RS image scene classification due to the availability of various spectral bands in the data.
-A central property of LRP and related relevancy-propagation-based methods such as Integrated Gradients [47] is relevance conservation: the total relevance assigned to the input for class $c$ equals the network output, $\sum_{j}r_{j}^{c}=f_{c}(\bm{x})$ [45].
-In natural images, discriminative information is typically concentrated within the object of interest.
-In contrast, RS scenes usually contain several LULC classes, so that a single class may occur in multiple, spatially disconnected regions; the discriminative information is therefore distributed rather than localized.
-Because the overall relevance budget is fixed, two classes that yield comparable output scores but occupy differently sized areas receive different relevance densities: the larger region is attributed a lower per-pixel relevance than the smaller region.
+Figure 3: LRP visualizations for an image from the DG-ML [46] dataset.
+a) Original image with two classes: Urban Land and Agricultural Land; b) Pixel-wise reference map (blue: Urban Land, orange: Agricultural Land); c) LRP explanation for Agricultural Land (normalized range: [0,1], unnormalized range: [0,0.02]); d) LRP explanation for Urban Land (normalized range: [0,1], unnormalized range: [0,0.18]).
+II.C.2 Limitations
+[…]
+A central property of LRP and related methods such as Integrated Gradients [52] is relevance conservation: the total input relevance equals the network output, $\sum_{j}r_{j}^{c}=f_{c}(\bm{x})$ [50].
+RS scenes usually contain several LULC classes, so that a single class may occur in multiple, spatially disconnected regions.
+The discriminative information is therefore distributed rather than localized.
+Because the overall relevance budget is fixed, two classes that yield comparable output scores but occupy differently sized areas receive different relevance densities.
+The larger region is attributed a lower per-pixel relevance than the smaller region.
 Fig.
-3 illustrates this issue using an image from the DG-ML dataset containing differently sized two classes: agricultural land ($c_{\text{major}}$) and urban land ($c_{\text{minor}}$).
-As the outputs of the model for both classes are similar $(f_{c_{\text{major}}}(\bm{x})\approxeq f_{c_{\text{minor}}}(\bm{x}))$ the maximum relevance to be propagated by LRP is similar.
-The higher number of relevant pixels for $c_{\text{major}}$ results in lower attributed relevance per pixel compared to $c_{\text{minor}}$.
-The $r^{c}_{j}$ for $c_{\text{major}}$ span values in the interval $[0,0.02]$, while the $r^{c}_{j}$ for $c_{\text{minor}}$ range within $[0,0.18]$.
-As attributions are typically normalized between 00 and $1$, this can produce attribution maps that are highly sensitive to outliers.
+3 shows this discrepancy for a DG-ML image containing agricultural land ($c_{\text{major}}$) and urban land ($c_{\text{minor}}$).
+Since $f_{c_{\text{major}}}(\bm{x})\approxeq f_{c_{\text{minor}}}(\bm{x})$, both propagate similar total relevance.
+However, the larger $c_{\text{major}}$ region yields lower per-pixel relevance ($[0,0.02]$) than $c_{\text{minor}}$ ($[0,0.18]$).
+As attributions are typically normalized to $[0,1]$, this discrepancy increases sensitivity to outliers.
 III Explanation metrics for RS image scene classification
@@
 We then introduce selected explanation metrics from different categories and discuss their suitability to estimate the quality of explanations for RS image scene classification.
-Explanation metrics offer a quantitative way to assess the quality of explanation methods.
-An explanation can be defined as a “presentation […] in human-understandable terms” [22].
-Because this definition centers on subjective human interpretation, it inherently lacks a single, absolute ground truth.
-Therefore, it is not possible to estimate an explanation error, as one would estimate the performance error of a DL model [24].
-Furthermore, explanation quality depends on more than the explanation method itself.
-When the decision-making process of a model is erroneous, even correctly derived explanations can appear inconsistent or counterintuitive, since explanations inherently reflect the internal logic of the model [48].
-To address these challenges, researchers have proposed metrics across several categories that capture distinct aspects of explanation quality.
-In our study, we consider five different categories of explanation metrics: Faithfulness, Robustness, Localization, Complexity and Randomization.
-Formally, the quality of an explanation is assessed using a metric $\Psi$, which evaluates the effectiveness of an explanation method:
+Explanation metrics quantify the quality of explanation methods.
+An explanation is commonly defined as a “presentation […] in human-understandable terms” [22].
+Since this definition depends on subjective interpretation, there is no absolute ground truth and thus no direct notion of explanation error [24].
+Furthermore, explanation quality also depends on the model itself: if the model’s reasoning is flawed, even faithful explanations may appear inconsistent or counterintuitive [53].
+To address these challenges, different metrics have been proposed across different categories, each capturing distinct aspects of explanation quality.
+We consider five categories of explanation metrics: Faithfulness, Robustness, Localization, Complexity, and Randomization.
+Formally, explanation quality is measured by a metric $\Psi$ evaluating the effectiveness of a method $\Phi$:
 $\Psi(\Phi,\bm{x},f,\hat{\bm{y}})=\hat{q}$, (18)
-where $\hat{q}$ is a scalar quality score. For simplicity, we omit explicit parameterization unless necessary.
-III A Faithfulness
-III A 1 Background
-The first category that we consider is faithfulness, which measures how well explanations reflect the actual decision-making process of the model.
-A faithful explanation accurately reflects why the model made a specific prediction.
-Metrics in this category assess this alignment by evaluating whether features identified as important, significantly influence the prediction of the model.
-Let us consider $\bm{x}$ as the input image and $V\subseteq[H]\times[W]$ a set of pixels to remove.
-We define the perturbation function $\mathcal{P}_{\mathbb{X}}(\bm{x},V)$ which replaces each feature in $V$ with a predefined baseline value.
-The general faithfulness category $\Psi_{\mathrm{F}}(\bm{x},V)$ is defined as:
-$\Psi_{\mathrm{F}}(\bm{x},V)=\bigl{\lvert}\,f(\bm{x})-f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr{)}\bigr{\rvert}.$ (22)
-Perturbations are commonly applied to the most relevant pixels first (MoRF) or the least relevant ones first (LeRF).
-To evaluate explanation methods for RS image scene classification, we consider two faithfulness metrics: i) Faithfulness Estimate (FE)[28], which measures the correlation between feature importance and prediction change, and ii) Iterative Removal Of Features (IROF)[29], which assesses prediction certainty decay when removing the most relevant segments.
-Faithfulness Estimate (FE)[28] quantifies the alignment between attribution magnitude and model sensitivity by measuring the correlation between the total attribution on the top-$K$ pixels and the resulting drop in prediction when those pixels are removed.
-Formally, we define the FE metric $\Psi_{\mathrm{FE}}$ as:
+where $\hat{q}$ is a scalar quality score. Explicit parameterization is omitted unless required.
+III.A Faithfulness
+III.A.1 Background
+Faithfulness measures how well explanations reflect the decision process of the model.
+A faithful explanation identifies features that truly influence predictions.
+Given input $\bm{x}$ and a set of pixels $V\subseteq[H]\times[W]$ to remove, the perturbation function $\mathcal{P}_{\mathbb{X}}(\bm{x},V)$ replaces the pixel in $V$ with a baseline value.
+Perturbations are applied either to the most relevant features first (MoRF), the least relevant first (LeRF), or randomly.
+We consider two faithfulness metrics: i) Faithfulness Estimate (FE) [28], which measures the correlation between feature importance and prediction change, and ii) Iterative Removal Of Features (IROF) [29], which evaluates prediction certainty decay when removing relevant segments.
+FE [28] measures the correlation between summed attributions on top-$K$ pixels and the prediction drop when those pixels are removed. Formally, we define the FE metric $\Psi_{\mathrm{FE}}$ as:
 $\Psi_{\mathrm{FE}}=\mathrm{corr}_{V:|V|=K}\Bigl(\sum_{(j,k)\in V}\Phi(\bm{x})_{j,k},\,f(\bm{x})-f\bigl(\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr)\Bigr).$ (19)
-Here, $\Phi(\bm{x})_{j,k}$ denotes the attributed relevancy at pixel $(j,k)$, and $V\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ is the set of the $K$ pixel indices with the largest attribution values.
-The function $\mathrm{corr}$ denotes the Pearson correlation coefficient.
-To compute $\Psi_{\mathrm{FE}}$, one first ranks all pixels by $\Phi(\bm{x})_{j,k}$, selects the top-$K$ indices $V$, and then correlates the summed attributions on $V$ with the change $f(\bm{x})-f\bigl{(}\mathcal{P}_{\mathbb{X}}(\bm{x},V)\bigr{)}$.
-A higher value of $\Psi_{\mathrm{FE}}$ indicates that the explanation method attributes the behavior of the model, i.e., i.e., removing features deemed important by the method results in a significant drop in prediction certainty.
-Iterative Removal Of Features (IROF)[29] measures how the prediction certainty changes when the most relevant features are removed but it considers the $K$ most relevant segments rather than pixels.
-Let us define $\{S_{m}\}_{m=1}^{M}$ as a partition of the image $\bm{x}$ into $M$ disjoint segments and sort them by their relevance, that:
-$\frac{1}{|S_{1}|}\sum_{(j,k)\in S_{1}}\Phi(\bm{x})_{j,k}\;\geq\;\cdots\;\geq\;\frac{1}{|S_{M}|}\sum_{(j,k)\in S_{M}}\Phi(\bm{x})_{j,k}.$ (24)
-For each $\kappa=0,1,\dots,K$, define the cumulative removal set $V_{\kappa}\;=\;\bigcup_{m=1}^{\kappa}S_{m},$ and the corresponding perturbed input $\mathcal{P}_{\mathbb{X}}(\bm{x},V_{\kappa}).$ Finally, the IROF score $\Psi_{\mathrm{IROF}}$ for $\bm{x}$ is obtained by normalizing each perturbed model output by the original confidence:
+Here, $\Phi(\bm{x})_{j,k}$ denotes the attributed relevancy at pixel $(j,k)$, and $V\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ is the set of the $K$ pixel indices with the largest attribution values and $\mathrm{corr}$ denotes the Pearson correlation coefficient.
+A high $\Psi_{\mathrm{FE}}$ value indicates that removing highly attributed pixels substantially lowers the prediction certainty.
+IROF [29] evaluates prediction certainty decay when removing the $K$ most relevant segments rather than pixels.
+Let $\{S_{m}\}_{m=1}^{M}$ be a partition of $\bm{x}$ into $M$ disjoint segments, which are ranked by their average attribution score.
+For $\kappa=0,\dots,K$, define the cumulative removal set $V_{\kappa}=\bigcup_{m=1}^{\kappa}S_{m}$ and perturbed input $\mathcal{P}_{\mathbb{X}}(\bm{x},V_{\kappa})$.
+The IROF metric $\Psi_{\mathrm{IROF}}$ is
 $\Psi_{\mathrm{IROF}}(\bm{x})=\frac{1}{K+1}\sum_{\kappa=0}^{K}\frac{f\bigl(\mathcal{P}_{\mathbb{X}}(\bm{x},V_{\kappa})\bigr)}{f(\bm{x})}.$ (20)
 A higher $\Psi_{\mathrm{IROF}}$ indicates that removing highly attributed segments causes a rapid drop in prediction confidence, and thus reflects a more faithful explanation.
-III A 2 Limitations
-Faithfulness metrics inherit several limitations from perturbation-based explanation methods, as both rely on input perturbations to evaluate changes in the prediction certainty of a model.
-Consequently, they exhibit similar issues to those discussed in Section II A 2.
-For example, like perturbation-based methods, most faithfulness metrics require a carefully chosen baseline to avoid distribution shifts that can affect the predictions of the model.
-In addition, the FE metric relies on a strong assumption: that removing a small number of high-attributed pixels will lead to a measurable change in model output.
+III.A.2 Limitations
+Faithfulness metrics inherit several limitations from perturbation-based explanation methods, as both rely on input perturbations to evaluate changes in the prediction certainty.
+Thus, they face similar issues to those in Section II.A.2, notably the need for carefully chosen baselines to avoid distribution shifts affecting model predictions.
+In addition, the FE metric relies on the assumption that removing a small number of highly attributed pixels will lead to a measurable change in model output.
 This assumption implicitly presumes that class-relevant information is concentrated in a small subset of pixels.
@@
 As a result, the Pearson correlation used by FE may become numerically unstable, since both numerator and denominator can approach zero, leading to undefined or misleading scores even for faithful explanations.
-In contrast to FE, some faithfulness metrics, such as IROF, evaluate the behavior of the model over a sequence of increasingly perturbed inputs.
-For these trajectory-based metrics, the order in which features are removed plays a critical role in shaping the evaluation outcome.
-To illustrate this, let us consider a scene in which a class $c_{\text{major}}$ occupies most of the image area, whereas a different class $c_{\text{minor}}$ covers only a small region.
-In such images, perturbing pixels belonging to $c_{\text{major}}$ do not significantly impact prediction certainty, as multiple pixels or regions contain similar predictive information.
-Consequently, metrics employing the MoRF strategy, which sequentially removes the most relevant pixels first, fail to capture meaningful changes in prediction certainty, as demonstrated by the slow decline of $f(\bm{x})$ in Fig.
+In contrast to FE, trajectory-based metrics like IROF assess model behavior across progressively perturbed inputs.
+Their outcome heavily depends on the feature removal order.
+Let us consider a scene where LULC class $c_{\text{major}}$ covers most of the image and $c_{\text{minor}}$ only a small region.
+Perturbing $c_{\text{major}}$ pixels has little effect due to redundancy, so removing most relevant pixels first yields a slow decline in $f(\bm{x})$, as visualized in Fig.
 4.
-In contrast, perturbations of the fewer, highly informative pixels associated with $c_{\text{minor}}$ lead to a rapid decrease in prediction certainty.
-This discrepancy results in biased metric outcomes, incorrectly suggesting that explanations for $c_{\text{minor}}$ are faithful, while explanations for $c_{\text{major}}$ appear less faithful.
-The LeRF strategy, which removes the least relevant pixels first, provides a better alternative to address this issue.
-By preserving the crucial predictive features of $c_{\text{minor}}$ until later perturbation steps, the LeRF strategy ensures a stable prediction certainty and yields a more balanced evaluation of explanation methods.
+In contrast, perturbing $c_{\text{minor}}$ pixels quickly lowers prediction certainty, biasing results toward faithful explanations for $c_{\text{minor}}$ and against $c_{\text{major}}$.
+The LeRF strategy, which removes the least relevant pixels first, mitigates this bias by preserving crucial features of $c_{\text{minor}}$ longer, stabilizing prediction certainty and producing more balanced evaluations.
 Figure 4: Comparison of MoRF and LeRF removal strategies for the target class $c$ (agricultural land, orange).
 Here, $K$ is the percentage of removed pixels and $f_{c}(\bm{x})$ represents the prediction certainty.
-Top row: a) Original image from the DG-ML dataset; b) Pixel-wise reference map.
+Top row: a) Original image from the DG-ML [46] dataset; b) Pixel-wise reference map.
 Agriculture land in orange; MoRF removal: c) $K=10\%$, $f_{c}(\bm{x})=0.98$; d) $K=50\%$, $f_{c}(\bm{x})=0.64$; e) $K=90\%$, $f_{c}(\bm{x})=0.31$.
 Bottom row: f) Grad-CAM attribution map for $c$; LeRF removal: g) $K=10\%$, $f_{c}(\bm{x})=0.99$; h) $K=50\%$, $f_{c}(\bm{x})=0.99$; i) $K=90\%$, $f_{c}(\bm{x})=0.72$.
-III B Robustness
-III B 1 Background
+III.B Robustness
+III.B.1 Background
 The second category we consider is robustness, which assesses the stability of explanations under small input perturbations.
-Similar, to adversarial attacks on DL models [49], explanations can also be fooled [50, 51].
-Thus, ensuring that the explanation is invariant to minimal perturbations in the input space is crucial for critical applications.
 We require small perturbations $\delta$ such that $\left\|\delta\right\|_{2}<\varepsilon$ and expect $\left\|\mathcal{P}_{\mathbb{X}}(\bm{x})-\bm{x}\right\|_{2}<\varepsilon$ for a small positive constant $\varepsilon$, assuming that the model output approximately stayed the same $f(\bm{x})\approx f\left(\mathcal{P}_{\mathbb{X}}(\bm{x})\right)$.
-The category $\Psi_{\mathrm{ROB}}$ can be defined as:
-$\Psi_{\mathrm{ROB}}(\Phi,\bm{x},\varepsilon)=\max_{\|\delta\|_{2}\leq\varepsilon}\bigl{\|}\,\Phi(\bm{x})\;-\;\Phi(\bm{x}+\delta)\bigr{\|}_{2}.$ (26)
 To evaluate the robustness of explanation methods for RS image scene classification, we use i) Average Sensitivity (AS) [30], which measures sensitivity to small perturbations using a Monte Carlo-based approximation and ii) Local Lipschitz Estimate (LLE) [28], which estimates Lipschitz continuity to assess explanation stability under minor input changes.
-Average Sensitivity (AS)[30] evaluates the sensitivity of the explanation method by perturbing samples using a Monte-Carlo-based approximation. The AS metric $\Psi_{\mathrm{AS}}$ is defined as:
+AS [30] evaluates the sensitivity of the explanation method by perturbing samples using a Monte-Carlo-based approximation. The AS metric $\Psi_{\mathrm{AS}}$ is defined as:
 $\Psi_{\mathrm{AS}}=\text{avg}_{\bm{x}+\delta\in\mathcal{N}_{\epsilon}(\bm{x})\leq\varepsilon}\left[\frac{\|\Phi(\bm{x})-\Phi(\bm{x}+\delta)\|}{\|\bm{x}\|}\right],$ (21)
 where $\varepsilon$ defines the radius of a discrete, finite-sample neighborhood around each input sample $\bm{x}$.
-This neighborhood denoted as $\mathcal{N}_{\epsilon}(\bm{x})$, includes all samples that are within a distance of $\varepsilon$ from $\bm{x}$.
+This neighborhood, denoted as $\mathcal{N}_{\epsilon}(\bm{x})$, includes all samples that are within a distance of $\varepsilon$ from $\bm{x}$.
 A lower score is indicative of more robustness.
-The Local Lipschitz Estimate (LLE)[28] estimates the Lipschitz continuity of the explanation, which measures how much the explanation changes with respect to the input under a minor perturbation, meaning the prediction did not change.
+The LLE [28] estimates the Lipschitz continuity of the explanation, which measures how much the explanation changes with respect to the input under a minor perturbation, meaning the prediction did not change.
 The LLE metric $\Psi_{\mathrm{LLE}}$ is defined as:
 $\Psi_{\mathrm{LLE}}=\max_{\bm{x}+\delta\in\mathcal{N}_{\epsilon}(\bm{x})\leq\epsilon}\left[\frac{|\Phi(\bm{x})-\Phi(\bm{x}+\delta)|_{2}}{|\bm{x}-(\bm{x}+\delta)|_{2}}\right],$ (22)
-where lower values indicate less change concerning the change in input, which is desirable.
-III B 2 Limitations
-A key hyper-parameter in this metric category is the noise type.
+where, $\delta$ is the perturbation vector with $\delta\in\mathcal{N}_{\epsilon}(\bm{x})\leq\epsilon$, and $\mathcal{N}\varepsilon(\bm{x})$ is the neighborhood of perturbed samples within radius $\varepsilon$.
+Here, lower values indicate less change concerning the change in input.
+III.B.2 Limitations
+A key hyperparameter in this metric category is the noise type.
 If the introduced noise does not naturally occur in the acquired images, it may generate OoD samples, reducing the reliability of the model and thereby affecting the validity of the explanations.
@@
 In contrast, for multispectral images, noise is typically independent of the data and generally additive in nature, such as additive white Gaussian noise [55, 56].
-III C Localization
-III C 1 Background
+III.C Localization
+III.C.1 Background
 The third category we consider is localization, which measures the spatial accuracy of explanations by comparing them to reference maps or bounding boxes.
 This comparison can be quantified using a similarity measure SIM between the generated explanation map $\Phi(\bm{x})$ and a pixel-level reference map $\bm{s}_{gt}$.
-The category $\Psi_{\mathrm{LOC}}$ is defined as:
-$\Psi_{\mathrm{LOC}}(\Phi,\bm{x},\bm{s}_{gt},\text{SIM})=\text{SIM}\left(\Phi(\bm{x}),\bm{s}_{gt}\right),$ (29)
 where SIM can be a measure such as Intersection over Union, Pearson correlation, or SSIM depending on the nature of the task and reference map.
-To evaluate the localization capabilities of explanation methods for RS image scene classification, we use i) Top-K Intersection (TKI)[32], which measures the fraction of top-ranked attributions within the pixel-level reference map, and ii) Relevance Rank Accuracy (RRA)[31], which assesses the concentration of high-relevance values within the reference map.
-Top-K Intersection (TKI), as proposed by Theiner et al. [32], measures the fraction of the $K$ highest ranked attributions that lie on the binary reference mask $\bm{s}_{gt}$. Let us define
-$I_{K}=\{\,i_{1},i_{2},\ldots,i_{K}\}\quad\text{and}\quad S_{gt}=\{\,p\mid\bm{s}_{gt}(p)=1\},$ (30)
-where $I_{K}$ is the set of the top-$K$ ranked pixel indices and $S_{gt}$ the set of true-positive pixels in the mask. The intersection of these sets is:
-$\bm{s}_{1-K}=I_{K}\;\cap\;S_{gt}.$ (31)
+We utilize i) Top-K Intersection (TKI) [32], which measures the fraction of top-ranked attributions within the pixel-level reference map, and ii) Relevance Rank Accuracy (RRA) [31], which assesses the concentration of high-relevance values within the reference map.
+The TKI metric [32] measures the fraction of the $K$ highest ranked attributions that lie on the binary reference mask $\bm{s}_{gt}$.
+We define $I_{K}=\{\,i_{1},i_{2},\ldots,i_{K}\}$ and $S_{gt}=\{\,p\mid\bm{s}_{gt}(p)=1\}$, where $I_{K}$ is the set of the top-$K$ ranked pixel indices and $S_{gt}$ the set of true-positive pixels in the mask.
+The intersection of these sets is $\bm{s}_{1-K}=I_{K}\;\cap\;S_{gt}$.
 The TKI metric $\Psi_{\mathrm{TKI}}$ is then defined as:
 $\Psi_{\mathrm{TKI}}=\frac{\bigl|\bm{s}_{1-K}\bigr|}{K}.$ (23)
-This measures how many of the $K$ most important features, according to $\bm{e}$, match the pixel‐level reference set $S_{gt}$.
-If all top-$K$ pixels are within the reference mask, then $\Psi_{\mathrm{TKI}}=1$, therefore a higher score is better.
-The RRA [31] measures the concentration of high-intensity relevance values within the pixel-level reference map.
-Let $|S_{\text{tot}}|$ be the size of the pixel-level reference map mask.
-Select the $K$ highest relevance values, according to $\Phi(\bm{x})$, count how many of these values lie within the actual pixel locations, and divide by $|S_{\text{tot}}|$.
-This can be expressed as:
-$P_{\text{top }K}=\left\{p_{1},p_{2},\ldots,p_{K}\mid R_{p_{1}}>R_{p_{2}}>\ldots>R_{p_{K}}\right\},$ (33)
-where $P_{\text{top }K}$ is the set of pixels with relevance values $R_{p_{1}},R_{p_{2}},\ldots,R_{p_{K}}$ sorted in decreasing order until the $K$-th pixel.
-The RRA metric $\Psi_{\mathrm{RRA}}$ is then computed as:
-$\Psi_{\mathrm{RRA}}=\frac{\left|P_{\text{top }K}\cap S_{\text{tot}}\right|}{|S_{\text{tot}}|}.$ (34)
-III C 2 Limitations
-In RS scene classification, label dependencies arise when the presence or absence of one label influences the likelihood of another, which is a common scenario in RS where certain objects (e.g., ships) typically appear in specific environments (e.g., sea).
-These dependencies are not considered by localization metrics.
-Furthermore, the reliability of localization metrics heavily depends on accurate and up-to-date pixel-level reference maps.
-Obtaining such annotations is costly and prone to errors.
-Inaccuracies or obsolescence in the reference maps can compromise the localization reliability, potentially leading to incorrect assessments.
-Additionally, most localization metrics become unreliable, when the target class spans a large area of the input image.
+If all top-$K$ pixels are within the reference mask $S_{gt}$, then $\Psi_{\mathrm{TKI}}=1$, therefore, a higher score is desired.
+The RRA metric [31] measures the concentration of high-intensity relevance values within the pixel-level reference map.
+Let $|S_{\text{tot}}|$ be the size of the reference mask.
+Let us assume that $K$ pixels with the highest relevance values are selected from $\Phi(\bm{x})$.
+We then count how many of these selected pixels lie within the reference pixels and divide this number by $|S_{\text{tot}}|$:
+$\Psi_{\mathrm{RRA}}=\frac{|P_{\text{top }K}\cap S_{\text{tot}}|}{|S_{\text{tot}}|}.$ (24)
+A higher $\Psi_{\mathrm{RRA}}$ indicates stronger localization, as more highly relevant pixels fall within the reference region.
+III.C.2 Limitations
+In RS scene classification, label dependencies arise when the presence of one label influences the likelihood of another, e.g., ships typically co-occur with sea.
+Localization metrics do not account for such dependencies.
+Their reliability also depends on accurate pixel-level reference maps, which are costly to obtain and prone to errors.
+Outdated or inaccurate annotations compromise the validity of these metrics.
+Additionally, most localization metrics become unreliable when the target class spans a large area of the input image.
 Let us consider the TKI metric in the scenario where we have a single label for the input image.
@@
 This phenomenon is well-documented for simpler localization metrics [57], such as the Pointing Game [58].
-However, in RS scene classification this also creates challenges for more sophisticated metrics, as in most RS scene classification tasks, as classes tend to span larger areas.
-III D Complexity
-III D 1 Background
+However, in RS scene classification, this also creates challenges for more sophisticated metrics, as in most RS scene classification tasks, as LULC classes tend to span larger areas.
+III.D Complexity
+III.D.1 Background
 The fourth category we consider is complexity, which quantifies the simplicity and interpretability of explanations, ensuring that they are useful for human understanding.
 These metrics evaluate properties such as sparsity, entropy, or other measures of complexity in the explanation map $\Phi(\bm{x})$.
-The category $\Psi_{\mathrm{CMP}}$ is defined as:
-$\Psi_{\mathrm{CMP}}(\Phi,f,\bm{x},C)=C\left(\Phi(\bm{x})\right),$ (35)
-where $C$ represents a complexity measure, chosen based on the context of interpretability.
-To evaluate the complexity of explanations for RS image scene classification, we use i) Sparseness (SP)[33] and ii) Complexity (CO)[34].
-Sparseness (SP)[33] is a method for evaluating the sparsity of explanations and is defined as the Gini index of the explanation.
+To evaluate the complexity of explanations for RS image scene classification, we use i) Sparseness (SP) [33]; and ii) Complexity (CO) [34].
+SP [33] is a method for evaluating the sparsity of explanations and is defined as the Gini index of the explanation.
 It is calculated by summing the product of the ranks of the input features and their attributions and dividing by the sum of the attributions.
@@
 A higher $\Psi_{\mathrm{SP}}$ indicates a sparser explanation.
-Complexity (CO)[34] is defined using the Shannon entropy calculation, which measures the amount of uncertainty or randomness in the attribution map.
+CO [34] is defined using the Shannon entropy calculation, which measures the amount of uncertainty or randomness in the attribution map.
 It is calculated by summing the product of the probabilities of the attributions and the logarithm of the probabilities of the attributions.
@@
 A uniformly distributed attribution would have the highest possible complexity score.
-Figure 5: Explanations for a DG-ML image labeled as Forest and their corresponding SP metric scores.
+Figure 5: Explanations for an image from the DG-ML dataset [46] labeled as Forest and their corresponding SP metric scores.
 a) Original image; b) Pixel-wise reference map (green: Forest); c) Occlusion explanation ($\Psi_{\mathrm{SP}}=0.08$); d) LRP explanation ($\Psi_{\mathrm{SP}}=0.47$).
-III D 2 Limitations
+III.D.2 Limitations
 Explanation metrics in the complexity category generally assume that an explanation is simpler if it attributes relevance to fewer features for a given prediction.
-This assumption originates from explaining predictions for natural images, where a class is often defined by a few unique features.
-There, only a small fraction of the input features contribute to the final prediction.
-However, classification of RS images often relies more on spectral and texture features, a larger set of repetitive features is relevant, which is overlooked by complexity metrics.
+This assumption originates from natural images, where classes are defined by a few distinctive features, and only a small fraction of pixels contribute to the prediction.
+However, as discussed before, in RS scenes LULC classes often cover large areas with repeated class-relevant features, so relevance is distributed across many pixels.
+Complexity metrics can therefore underestimate explanation quality.
 A key limitation of the Gini-based SP metric $\Psi_{\mathrm{SP}}$ is its tendency to penalize explanations with uniformly distributed attributions.
@@
 This limitation is illustrated in Fig.
-5, which shows two explanations for a scene labeled as forest, a class that spans nearly the entire image.
+5, which shows two explanations for an image labeled as forest, a class that spans nearly the entire image.
 The attribution map from the Occlusion method (Fig.
@@
 Although most of the input pixels are attributed as relevant, the map remains easy to interpret.
-By contrast, the LRP map (Fig.
+On the contrary, the LRP map (Fig.
 5(d)) concentrates relevance on a small, localized area, producing a higher sparsity score ($\Psi_{\mathrm{SP}}=0.47$).
-Here the SP metric favors a compact but misleading attribution, mistaking spatial concentration for interpretive simplicity.
-III E Randomization
-III E 1 Background
+Here, the SP metric favors a compact but misleading attribution, mistaking spatial concentration for interpretive simplicity.
+III.E Randomization
+III.E.1 Background
 The fifth category we consider is randomization, which tests the reliability of explanations by introducing randomness into model parameters or prediction targets.
 These metrics evaluate the consistency of explanations under such perturbations to ensure that explanations are not artifacts of specific parameter settings or configurations.
-The category $\Psi_{\mathrm{RAND}}$ is defined as:
-$\Psi_{\mathrm{RAND}}(\Phi,f,\bm{x},\mathcal{R},\text{SIM})=\text{SIM}\left(\Phi(\bm{x}),\Phi(\bm{x},f,\mathcal{R})\right),$ (39)
-where $\mathcal{R}$ represents a randomization operation (e.g., model parameter shuffling or target randomization) and SIM measures the similarity between the original explanation $\Phi(\bm{x},f)$ and the explanation generated under the randomized condition $\Phi(\bm{x},f,\mathcal{R})$.
-To evaluate the randomization robustness of explanation methods for RS image scene classification, we use i) Efficient Model Parameter Randomization Test (MPRT)[57], which measures explanation complexity under model parameter randomization and ii) Random Logit (RL)[36], which evaluates stability by comparing explanations for target and non-target classes using Structural Similarity Index Measure (SSIM).
-Efficient Model Parameter Randomization Test (MPRT)[57] measures the relative rise in the complexity of the explanation $\Phi\bigl{(}\bm{x},f_{\mathcal{R}}\bigr{)}$ from a fully randomized model $f_{\mathcal{R}}$ for image $\bm{x}$.
+Here, we use i) Efficient Model Parameter Randomization Test (MPRT) [59], which measures explanation complexity under model parameter randomization, and ii) Random Logit (RL) [36], which evaluates stability by comparing explanations for target and non-target classes using Structural Similarity Index Measure (SSIM).
+MPRT [59] measures the relative increase in the complexity of the explanation $\Phi\bigl(\bm{x},f_{\mathcal{R}}\bigr)$ from a fully randomized model $f_{\mathcal{R}}$ for image $\bm{x}$.
 The MPRT metric $\Psi_{MPRT}$ is defined as:
@@
 To define $\xi$, a histogram‐entropy measure grounded in Shannon entropy is employed.
-Specifically, the vectorized attributions $\bm{e}=\{\,a_{j,k}\}_{j,k}$ are placed into $B$ distinct bins; for the $d$-th bin, the frequency $\eta_{d}$ is computed and the corresponding normalized probability $\pi_{d}$ is
+Specifically, the vectorized attributions $\bm{e}=\{\,a_{j,k}\}_{j,k}$ are placed into $B$ distinct bins.
+For the $d$-th bin, the frequency $\eta_{d}$ is computed and the corresponding normalized probability $\pi_{d}$ is
 $\xi(\bm{e})=-\sum_{d=1}^{B}\pi_{d}\,\log\!\bigl(\pi_{d}\bigr),\quad\text{where}\quad\pi_{d}=\frac{\eta_{d}}{\sum_{d^{\prime}=1}^{B}\eta_{d^{\prime}}}\,.$ (29)
-The Random Logit (RL)metric [36] is defined as the SSIM over the explanation of the reference label and an explanation of non-target class $\bm{y}^{\prime}$.
-The RS metric $\Psi_{\mathrm{RL}}$ is defined as:
+The RL metric [36] is defined as the SSIM over the explanation of the reference label and an explanation of the non-target class $\bm{y}^{\prime}$. The RS metric $\Psi_{\mathrm{RL}}$ is defined as:
 $\Psi_{\mathrm{RL}}=\operatorname{SSIM}\left(\Phi(\bm{x},f,\hat{\bm{y}}),\Phi\left(\bm{x},f,\bm{y}^{\prime}\right)\right),$ (30)
-where $\Phi(\bm{x},f,\hat{\bm{y}})$ is the explanation generated for the prediction $\hat{\bm{y}}$ and $\Phi\left(\bm{x},f,\bm{y}^{\prime}\right)$ is the explanation generated for a non-target class $\bm{y}^{\prime}$.
-Lower values indicate that the explanations are not correlated, which is desirable.
-III E 2 Limitations
+[…]
+III.E.2 Limitations
 Due to their minimal sensitivity to variations in input data characteristics, randomization metrics are also robust for RS image classification.
@@
 IV Dataset Description and Experimental Design
-IV A Dataset Description
+IV.A Dataset Description
 The experiments are conducted using three datasets: 1) BigEarthNet-S2 v2.0 (BEN-v2.0) [60], 2) DeepGlobe Multi-Label (DG-ML) [46, 61], and 3) Five Billion Pixel Multi-Label (FBP-ML) [62].
@@
 per Image
-BEN-v2.0 $250.249$ $19$ $2.95$ $10$ 120$\times$120 (10 m) Thematic Product
-DG-ML $18.185$ $6$ $1.71$ $3$ 120$\times$120 (0.5 m) Manually Annotated
-FBP-ML $109.200$ $24$ $2.61$ $4$ 256$\times$256 (4 m) Manually Annotated
-BEN-v2.0 [58] is a multi-label dataset which extends the BigEarthNet archive [62] and covers ten countries in Europe.
-We use the Sentinel-2 imagery with 12 bands and 10 m spatial resolution.
-Each $1200$ m×$\times 1200$ m tile is subdivided into $120\times 120$-pixel patches.
-Class annotations are taken from the CORINE Land-Cover Map inventory 2018, assigning subsets of 19 LULC classes to 549 488 patches.
-We adopt the official train–validation–test split defined in [58].
+BEN-v2.0 [60] $250.249$ 1919 $2.95$ $10$ $\times120$ (10 m) Thematic Product
+DG-ML [46] $18.185$ 66 $1.71$ 33 $\times120$ (0.5 m) Manually Annotated
+FBP-ML [62] $109.200$ $24$ $2.61$ 44 $\times256$ (4 m) Manually Annotated
+BigEarthNet-S2 v2.0 (denoted as BEN-v2.0 hereafter) [60] is a multi-label dataset, including 549 488 Sentinel-2 images acquired over ten countries in Europe.
+We used all the bands of the Sentinel-2 images, and thus applied cubic interpolation to 20-m and 60-m bands.
+Each image is associated with one or more class labels (i.e., multi-labels) based on the 19 class nomenclature [63].
+In the experiments, we use the official train-validation-test split defined in [60].
 DG-ML is a multi-label dataset, which Burgert et al.
-[60] derived from the original DeepGlobe Land-Cover Classification Challenge [59].
+[61] derived from the original DeepGlobe Land-Cover Classification Challenge dataset [46].
 The original DeepGlobe corpus comprises 1,949 manually annotated RGB tiles of size $2448\times 2448$ pixels at 0.5 m spatial resolution, collected over Thailand, Indonesia, and India.
-The LULC classes are urban land, agricultural land, range land, forest, water, barren land, and unknown.
-Each tile was divided into a grid of $120\times 120$-pixel patches, and multi-labels were assigned from the corresponding pixel-level reference maps, while patches containing the class “unknown” were discarded.
-Only 20 % of single-class patches and all multi-class patches are retained, yielding 30 443 samples split into training (60%), validation (20%) and test (20%) subsets.
+Each tile is divided into 400 images, each of which is a section of $120\times 120$ pixel images.
+Multi-labels were assigned from the corresponding pixel-level reference maps, while images containing the class “unknown” were discarded.
+Only 20% of single-class images and all multi-class images are retained, yielding 30 443 samples split into training (60%), validation (20%), and test (20%) subsets.
 FBP-ML is a multi-label dataset derived from the Five Billion Pixels (FBP) collection [62], comprising 150 Gaofen-2 RGB–NIR scenes (4 m spatial resolution) that together cover roughly 50 000 km2 in China.
-Following the procedure used for DG-ML, each scene is partitioned into a grid of $256\times 256$-pixel patches and labeled with a 24-class scheme that distinguishes artificial, agricultural, and natural surfaces.
-The final dataset consists of 87 360 training samples (80 %), 10 920 validation samples (10 %), and 10 920 test samples (10 %).
-IV B Evaluation Protocol
+Following the procedure used for DG-ML, each tile is partitioned into an area of $256\times 256$-pixel images and labeled with a 24-class scheme that distinguishes artificial, agricultural, and natural surfaces.
+The final dataset consists of a training set with 87 360 samples (80%), a validation set with 10 920 samples (10%), and a test set with 10 920 samples (10%).
+IV.B Evaluation Protocol
 In the experiments, we evaluate explanation methods and metrics for feature attribution in RS image scene classification using the pipeline illustrated in Fig.
@@
 Then, the reliability of these metrics is assessed by using the MetaQuantus [24] framework.
-The MetaQuantus framework [24] considers that the quality of an explanation cannot be objectively measured due to the lack of ground truth.
-It tests the explanation metrics against two failure modes: Noise Resilience (NR)of a metric, which is the resilience against minor perturbations, and Adversary Reactivity (AR), which is the reactivity of the metrics to disruptive perturbations.
-A perturbation is considered minor when the predicted class label does not change and disruptive when the predicted class label changes.
+As the quality of an explanation cannot be objectively measured due to the lack of ground truth [24], the explanation metrics are tested against two failure modes: Noise Resilience (NR) of a metric, which is the resilience against minor perturbations, and Adversary Reactivity (AR), which is the reactivity of the metrics to disruptive perturbations.
+A perturbation is considered minor when the predicted class label does not change, and disruptive when the predicted class label changes.
 To test against these failure modes, two consistency criteria Intra-Consistency (IAC) and Inter-Consistency (IEC) are utilized.
@@
 The entries $U_{i,j}^{t}$ capture how consistently the ranking of the metrics agrees between unperturbed $\bar{\mathbf{Q}}$ and perturbed $\overline{\mathbf{Q}}^{\prime}$.
-For minor perturbations ($t=M$) the ordering of the $L$ methods for each sample must remain identical, whereas for disruptive perturbations ($t=D$) the unperturbed score must always ranked higher than its perturbed counterpart.
+For minor perturbations ($t=M$) the ordering of the $L$ methods for each sample must remain identical, whereas for disruptive perturbations ($t=D$), the unperturbed score must always be ranked higher than its perturbed counterpart.
 The Meta-Consistency (MC) score combines both intra- and inter-consistency criteria across different failure modes as:
@@
 with the all-ones vector representing an ideal explanation metric that achieves a value of 1 on all criteria.
-A metric with an ${MC}$ score close to 1 indicates strong performance across all criteria tested with higher values reflecting greater overall robustness and reliability.
-The MC score is averaged over two types of perturbation: model perturbation and input perturbation.
+A metric with an $\mathbf{MC}$ score close to 1 indicates strong performance across all criteria tested, with higher values reflecting greater overall robustness and reliability.
+The $\mathbf{MC}$ score is averaged over two types of perturbation: model perturbation and input perturbation.
 In model perturbation, random noise is added to the parameters, whereas in input perturbation, it is added to the input data.
-IV C Experimental Setup
+While the protocol provides a systematic assessment of explanation metrics, it is computationally demanding as its runtime depends on the efficiency of the employed explanation methods and metrics.
+In particular, perturbation-based methods and metrics require multiple forward passes and repeated perturbations, which significantly increases computation time compared to closed-form or gradient-based approaches.
+This limitation is not inherent to the evaluation protocol but reflects the methodological design of certain metrics.
+A detailed comparison of computational efficiency across methods and metrics is provided in subsection V.C.
+IV.C Experimental Setup
 We have systematically evaluated ten explanation metrics across five widely used feature attribution methods on multi-label RS datasets (BEN-v2.0, DG-ML, and FBP-ML).
@@
 We evaluated the quality of the explanation using the considered metrics and evaluated their reliability with MetaQuantus [24] on 512 randomly drawn samples, a sample size demonstrated to be sufficient in previous work [24].
-The hyperparameter were set as in the original publication ($k=5$, three iterations).
+The hyperparameters were set as in the original publication ($k=5$, three iterations, which were shown in [24] to be effective across different datasets.
 Gaussian noise was chosen for the input perturbation to preserve physical consistency, in line with the recommendations in [64].
@@
 For a multi-label case with a label set $\{1,…,K\}$, we apply the same method separately to each label $c\in\{1,…,K\}$ to obtain a feature attribution map for each class $c$.
-Furthermore, as common in the literature, we summed the feature attributions over the channels.
+Furthermore, as is common in the literature, we summed the feature attributions over the channels.
 The hyperparameters for the explanation methods were configured as follows.
-For DeepLIFT, the perturbation baseline value was set to 00.
-For GradCAM, the last convolutional layer was analyzed and the explanations were upscaled by bilinear interpolation.
+For DeepLIFT, the perturbation baseline value was set to 0.
+For GradCAM, the last convolutional layer was analyzed, and the explanations were upscaled by bilinear interpolation.
 The LIME method employed a Euclidean exponential kernel (kernel width 500) as a similarity measure, used linear regression as an interpretable model, and generated 15 superpixels using SLIC.
-For LRP, following [45], the first third of the network was initialized with the LRP-$\gamma$ rule, the second third with the LRP-$\epsilon$ rule and the final third with the LRP-00 rule.
-Lastly, the Occlusion method employed a sliding window of size $(50,50)$ with strides of $(10,10)$ across the channels and the perturbation baseline was set to 00.
+For LRP, following [50], the first third of the network was initialized with the LRP-$\gamma$ rule, the second third with the LRP-$\epsilon$ rule and the final third with the LRP-0 rule.
+Lastly, the Occlusion method employed a sliding window of size $(50,50)$ with strides of $(10,10)$ across the channels, and the perturbation baseline was set to 0.
 The quality of the explanations was assessed using 10 explanation metrics implemented via the Quantus framework [25] with default hyperparameters.
@@
 The explanation metrics were computed over 1024 randomly drawn test samples from each dataset.
+Both explanation methods and metrics were used with their standard parameters to evaluate the configurations that are most commonly used in RS.
 The explanation methods are implemented using the Captum framework [66] with modifications to Captum, Quantus, and MetaQuantus to support multi-label image scene classification.
+Figure 6: Mean $\mathbf{MC}$ scores for the considered explanation metrics: a) scores associated to all considered RS datasets (DG-ML [46], FBP-ML [62] and BEN-v2.0 [60]); b) scores associated to each individual dataset (DG-ML [46], FBP-ML [62] and BEN-v2.0 [60]).
+Metrics are categorized into five color-coded groups: i) Complexity (brown), ii) Faithfulness (blue), iii) Localization (orange), iv) Randomization (green), and v) Robustness (red).
 V Experimental Results
@@
 This ordering ensures that the selection of suitable explanation methods is grounded in metrics whose reliability has been rigorously validated.
-V A Evaluation of the considered explanation metrics
-Figure 6: Mean MC scores for the considered explanation metrics: a) scores associated to all RS datasets (DG-ML, FBP-ML and BEN-v2.0); b) scores associated to each individual dataset (DG-ML, FBP-ML, BEN-v2.0).
-Metrics are categorized into five color-coded groups: Complexity (brown), Faithfulness (blue), Localization (orange), Randomization (green), and Robustness (red).
+V.A Evaluation of the considered explanation metrics
 In Fig.
@@
 In contrast, Fig.
-6(b) decomposes these results by dataset: for each explanation metric, three separate bars show its mean MC score on DG-ML, FBP-ML and BEN-v2.0 individually.
+6(b) decomposes these results by dataset.
+For each explanation metric, three separate bars show its mean MC score on DG-ML, FBP-ML, and BEN-v2.0 individually.
 Across both figures, the magnitude of the metrics remains stable, indicating that metric reliability is robust whether computed on the combined dataset or on each dataset in isolation.
 From Fig.
-6(a) one can see that the metrics from the robustness category are the most reliable for RS image scene classification.
+6(a), one can see that the metrics from the robustness category are the most reliable for RS image scene classification.
 In particular, LLE achieves the highest mean MC score of $0.79$, while AS attains the second-highest score of $0.68$.
-This is because the metrics in this category have fewer limitations compared to other categories and are therefore particularly well-suited for RS image scene classification (see Section III B 2).
+This is because the metrics in this category have fewer limitations compared to other categories and are therefore particularly well-suited for RS image scene classification (see Section III.B.2).
 By observing the figure, one can also see that the localization metrics, RRA and TKI, perform worst with mean MC scores of $0.52$ and $0.51$, respectively.
-The primary cause of the low performance of localization metrics arises from their methodological constraint: They are unreliable when there is a single LULC class present in $\bm{x}$.
+The primary cause of the low performance of localization metrics arises from their methodological constraint, which is that they are unreliable when there is a single LULC class present in $\bm{x}$.
 On a closer look, the evaluation of the input perturbation test on these metrics (TKI in Fig.
@@
 Specifically, the input perturbation test yields similar trends across the four evaluation scores that contribute to the MC score (see Eq.
-46).
+34).
 Both TKI and RRA exhibit high values for $\text{IAC}_{\text{NR}}$ and relatively high $\text{IEC}_{\text{NR}}$ values.
 On the other hand, both metrics show low values for $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$.
-Figure 7: Results of the evaluation of the input test on the DG-ML dataset for the localization metrics: a) Top-K Intersection and b) Relevance Rank Accuracy.
+Figure 7: Results of the evaluation protocol on the DG-ML dataset [46] for the localization metrics: a) Top-K Intersection and b) Relevance Rank Accuracy.
 The NR scores indicate consistency over minor perturbations, while the AR scores reflect changes in response to disruptive perturbations.
-The pattern observed in these scores responds to the methodological flaws of the localization metrics described in Section III C 2.
+The pattern observed in these scores responds to the methodological flaws of the localization metrics described in Section III.C.2.
 Specifically, when a single LULC class spans the entire image, the explanation coincides exactly with the reference map.
 Therefore, the metric always yields an output value of $1$.
-Since the $\text{IAC}_{\text{ER}}$ and $\text{IEC}_{\text{ER}}$ scores measure the metrics consistency over minor perturbations, both scores will consistently assign the highest score of 1.
+Since the $\text{IAC}_{\text{ER}}$ and $\text{IEC}_{\text{ER}}$ scores measure the metrics consistency over minor perturbations, both scores consistently assign the highest score of 1.
 However, as the output is constant, disruptive perturbations cannot change the output of the metric, leading to low $\text{IAC}_{\text{AR}}$ and $\text{IEC}_{\text{AR}}$ values.
@@
 Furthermore, it is visible in Fig.
-6(a) that the randomization metrics also perform comparatively well with the second highest average scores.
+6(a) that the randomization metrics also perform comparatively well with the second-highest average scores.
 This is in accordance with the assumption that the metrics of this category perform independently of the data.
 From the figure, one can see that the MPRT metric achieves a mean MC score of $0.68$, while the RL metric attains a score of $0.60$.
-Figure 8: Results of the evaluation of the input test on the DG-ML dataset for the complexity metrics: a) Complexity; and b) Sparseness.
+Figure 8: Results of the evaluation protocol on the DG-ML dataset [46] for the complexity metrics: a) Complexity; and b) Sparseness.
 The mean MC values for the metrics CO and SP in the complexity category are $0.63$ and $0.57$, as visualized in Fig.
 6(a).
-Similar to the localization metrics, the complexity metrics also struggle with the characteristic of RS image data in the case that single LULC classes can span large parts of an image, yet only a small region provides the discriminative information necessary for accurate classification.
+Similar to the localization metrics, the complexity metrics also struggle with the characteristic of RS image data.
+Single LULC classes can span large parts of an image, yet only a small region provides the discriminative information necessary for accurate classification.
 In this case, the metrics tend to attribute higher complexity to the explanation.
@@
 8(b).
-Figure 9: Results of the evaluation of the input test on the BEN-v2.0 dataset for the faithfulness metrics: a) Faithfulness Estimate; and b) Iterative Removal Of Features.
+Figure 9: Results of the evaluation of the evaluation protocol on the BEN-v2.0 dataset [60] for the faithfulness metrics: a) Faithfulness Estimate; and b) Iterative Removal Of Features.
 As shown in Fig.
@@
 9(a) and Fig.
-9(b) present the detailed evaluation results for the BEN-v2.0 dataset, which we consider as the perturbation baseline is expected to have a stronger impact for multi-spectral RS images.
-Similar to the localization metrics, the $\text{IEC}_{\text{AR}}$ score is notably low, with values of 0.28 for FE and 0.01 for IROF.
+9(b) present the detailed evaluation results for the BEN-v2.0 dataset, which we consider as the perturbation baseline is expected to have a stronger impact for multispectral RS images.
+Similar to the localization metrics, the $\text{IEC}_{\text{AR}}$ score is notably low with values of 0.28 for FE and 0.01 for IROF.
 This aligns with methodological findings that faithfulness metrics face similar challenges as localization metrics when an image contains redundant image features for classes that span large parts of the input image.
-In such cases, the prediction certainty is not expected to decrease significantly when important features are removed (see Section III A 2).
+In such cases, the prediction certainty is not expected to decrease significantly when important features are removed (see Section III.A.2).
 A low $\text{IEC}_{\text{AR}}$ indicates that the metric scores for the perturbed input are consistently lower than for the unperturbed input.
 However, when the prediction of the model remains unchanged, faithfulness metrics still assign a high score, keeping the score on the unperturbed input high and consequently leading to low $\text{IEC}_{\text{AR}}$ values.
-Conversely, the $\text{IAC}_{\text{AR}}$ scores for FE and IROF are high, with values of $0.82$ and $1.0$, respectively.
+On the contrary, the $\text{IAC}_{\text{AR}}$ scores for FE and IROF are high, with values of $0.82$ and $1.0$, respectively.
 This suggests that the distribution of scores remains unchanged even under disruptive perturbations.
-One possible explanation is a distribution shift induced by OoD samples that occurs due to the perturbation, which prevents the faithfulness scores from effectively capturing the expected changes.
-Figure 10: Scores of the IROF metric on the BEN-v2.0 dataset under four perturbation configurations: (i) mean-based LeRF, (ii) mean-based MoRF, (iii) black-based MoRF, and (iv) uniform-based MoRF.
-To further investigate how perturbation baselines and removal strategies influence the reliability of faithfulness metrics on multispectral images, we conduct an ablation study using the IROF metric with four configurations: mean-based LeRF, mean-based MoRF, black-based MoRF, and uniform-based MoRF (Fig.10).
+One possible explanation is a distribution shift caused by OoD samples introduced through the perturbation, which prevents the faithfulness scores from effectively capturing the expected changes.
+Figure 10: MC Scores of the IROF metric on the BEN-v2.0 dataset [60] under four perturbation configurations: (i) mean-based LeRF, (ii) mean-based MoRF, (iii) black-based MoRF, and (iv) uniform-based MoRF.
+To further investigate how perturbation baselines and removal strategies influence the reliability of faithfulness metrics on multispectral images, we conduct an ablation study using the IROF metric with four configurations: i) mean-based LeRF; ii) mean-based MoRF; iii) black-based MoRF; and iv) uniform-based MoRF (Fig.10).
 The results show that combining the mean baseline with LeRF yields the highest reliability (0.75), followed by mean-based MoRF (0.64), black-based MoRF (0.61), and uniform-based MoRF (0.58).
-As discussed in Section III A 2, removing the most relevant pixels first can fail to capture meaningful changes in prediction certainty, particularly when large image regions contain similar predictive information.
+As discussed in Section III.A.2, removing the most relevant pixels first can fail to capture meaningful changes in prediction certainty, particularly when large image regions contain similar predictive information.
 In contrast, the LeRF approach preserves critical features until later perturbation steps, leading to a more balanced evaluation.
 Moreover, employing a mean-value baseline reduces the likelihood of generating OoD samples, further enhancing the reliability of the faithfulness scores.
-V B Evaluation of the considered feature attribution methods
-TABLE II: Average metric scores obtained by the considered explanation methods for BEN-v2.0
+V.B Evaluation of the considered feature attribution methods
+TABLE II: Average metric scores obtained by the considered explanation methods for BEN-v2.0 [60].
 Faithfulness $\uparrow$ Robustness $\uparrow$ Localization $\uparrow$ Complexity $\uparrow$ Randomization $\uparrow$
-Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
-DeepLIFT 0.66 0.59 0.49 1.00 0.49 0.48 0.60 0.30 0.40 0.49
-GradCAM 0.73 0.56 0.88 0.99 0.50 0.43 0.46 0.47 0.81 0.67
-LIME 0.61 0.54 0.50 0.05 0.49 0.44 0.58 0.46 0.66 0.26
-LRP 0.61 0.51 0.71 0.99 0.48 0.45 0.55 0.20 0.23 0.25
-Occlusion 0.64 0.48 0.87 0.71 0.49 0.44 0.35 0.52 0.74 0.42
+Explanation Method FE [28] IROF [29] AS [30] LLE [28] TKI [32] RRA [31] SP [33] CO [34] RL [36] MPRT [59]
+DeepLIFT [12] 0.66 0.59 0.49 1.00 0.49 0.48 0.60 0.30 0.40 0.49
+GradCAM [13] 0.73 0.56 0.88 0.99 0.50 0.43 0.46 0.47 0.81 0.67
+LIME [10] 0.61 0.54 0.50 0.05 0.49 0.44 0.58 0.46 0.66 0.26
+LRP [11] 0.61 0.51 0.71 0.99 0.48 0.45 0.55 0.20 0.23 0.25
+Occlusion [9] 0.64 0.48 0.87 0.71 0.49 0.44 0.35 0.52 0.74 0.42
 Random baseline 0.39 0.44 0.85 0.60 0.47 0.42 0.07 0.00 0.96 0.26
-TABLE III: Average metric scores obtained by the considered explanation methods for DG-ML
+TABLE III: Average metric scores obtained by the considered explanation methods for DG-ML [46].
 Faithfulness $\uparrow$ Robustness $\uparrow$ Localization $\uparrow$ Complexity $\uparrow$ Randomization $\uparrow$
-Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
-DeepLIFT 0.67 0.65 0.41 1.00 0.62 0.45 0.71 0.48 0.46 0.66
-GradCAM 0.65 0.65 0.68 0.96 0.74 0.48 0.43 0.42 0.83 0.49
-LIME 0.53 0.66 0.66 0.56 0.64 0.44 0.52 0.50 0.72 0.23
-LRP 0.69 0.66 0.61 0.99 0.64 0.47 0.54 0.26 0.22 0.42
-Occlusion 0.75 0.68 0.69 0.43 0.69 0.47 0.31 0.40 0.82 0.34
+Explanation Method FE [28] IROF [29] AS [30] LLE [28] TKI [32] RRA [31] SP [33] CO [34] RL [36] MPRT [59]
+DeepLIFT [12] 0.67 0.65 0.41 1.00 0.62 0.45 0.71 0.48 0.46 0.66
+GradCAM [13] 0.65 0.65 0.68 0.96 0.74 0.48 0.43 0.42 0.83 0.49
+LIME [10] 0.53 0.66 0.66 0.56 0.64 0.44 0.52 0.50 0.72 0.23
+LRP [11] 0.69 0.66 0.61 0.99 0.64 0.47 0.54 0.26 0.22 0.42
+Occlusion [9] 0.75 0.68 0.69 0.43 0.69 0.47 0.31 0.40 0.82 0.34
 Random baseline 0.44 0.57 0.84 0.78 0.59 0.37 0.19 0.03 0.97 0.36
-TABLE IV: Average metric scores obtained by the considered explanation methods for FBP-ML
+TABLE IV: Average metric scores obtained by the considered explanation methods for FBP-ML [62].
 Faithfulness $\uparrow$ Robustness $\uparrow$ Localization $\uparrow$ Complexity $\uparrow$ Randomization $\uparrow$
-Explanation Method FE IROF AS LLE TKI RRA SP CO RL MPRT
-DeepLIFT 0.75 0.59 0.45 1.00 0.46 0.49 0.70 0.43 0.29 0.69
-GradCAM 0.75 0.62 0.75 0.99 0.48 0.50 0.57 0.85 0.71 0.46
-LIME 0.65 0.55 0.56 0.41 0.45 0.45 0.55 0.51 0.67 0.16
-LRP 0.69 0.58 0.53 0.99 0.43 0.47 0.65 0.28 0.29 0.26
-Occlusion 0.81 0.58 0.69 0.81 0.46 0.48 0.50 0.48 0.73 0.25
+Explanation Method FE [28] IROF [29] AS [30] LLE [28] TKI [32] RRA [31] SP [33] CO [34] RL [36] MPRT [59]
+DeepLIFT [12] 0.75 0.59 0.45 1.00 0.46 0.49 0.70 0.43 0.29 0.69
+GradCAM [13] 0.75 0.62 0.75 0.99 0.48 0.50 0.57 0.85 0.71 0.46
+LIME [10] 0.65 0.55 0.56 0.41 0.45 0.45 0.55 0.51 0.67 0.16
+LRP [11] 0.69 0.58 0.53 0.99 0.43 0.47 0.65 0.28 0.29 0.26
+Occlusion [9] 0.81 0.58 0.69 0.81 0.46 0.48 0.50 0.48 0.73 0.25
 Random baseline 0.37 0.39 0.85 0.45 0.45 0.41 0.06 0.00 0.98 0.17
@@
 Randomization is quantified using the RL and MPRT metrics.
-The random baseline achieves the highest RL score across all datasets, with an average performance of 0.97, aligning with the findings of [66], which demonstrate that randomized attribution maps can achieve high scores in top-down randomization metrics.
+The random baseline achieves the highest RL score across all datasets, with an average performance of 0.97.
+This aligns with the findings of [67], which demonstrate that randomized attribution maps can achieve high scores in top-down randomization metrics.
 While this outcome exposes a limitation of the RL metric, it remains valuable in identifying explanation methods that are largely independent of the model and therefore unreliable.
@@
 Fig.
-12(a) displays the overall average performance of the evaluated explanation methods across all considered metrics.
-The results show that GradCAM achieves the highest score ($0.65$), followed by DeepLIFT ($0.58$) and Occlusion ($0.57$).
-LRP and LIME achieve slightly lower scores of $0.52$ and $0.50$, respectively, while the Random Baseline performs worst at $0.46$.
-In contrast, Fig.
-12(b) presents the weighted average scores of the most reliable metrics.
-Here, GradCAM again performs best ($0.66$), but DeepLIFT achieves a higher relative score ($0.63$), surpassing Occlusion ($0.52$).
-Notably, LRP improves in ranking, reaching a score of $0.53$, while LIME and the Random Baseline exhibit a more pronounced decline, with scores of $0.43$ and $0.37$, respectively.
-Comparing both figures, GradCAM consistently outperforms other methods, maintaining the highest rank across both evaluation approaches.
-However, the usage of more reliable metrics in Fig.
-12(b) results in an increased performance gap between methods, with DeepLIFT improving its relative standing and LIME and the Random Baseline exhibiting a sharper decline.
-This shows that using more reliable metrics leads to a more accurate evaluation and confirms the robustness of GradCAM.
-The larger reduction in Random Baseline performance indicates that the most reliable metrics for RS image scene classification effectively penalize random explanations.
-Figure 11: Comparison of metric averages across datasets. Each metric is color-coded by category.
+12(a) and Fig.
+12(b) show a comparison of the overall and weighted average performances of the evaluated explanation methods across all considered metrics.
+In both evaluations, GradCAM consistently achieves the highest scores.
+When weighting by the most reliable metrics, the relative differences between methods become more pronounced.
+In detail, DeepLIFT improves its standing, while LIME and the Random Baseline show a substantial decline.
+This indicates that reliable metrics provide a sharper separation between strong and weak methods, effectively penalizing random explanations and demonstrating the reliability of GradCAM.
+Figure 11: Average performance of explanation metrics for each considered dataset, grouped by evaluation category.
+Each bar corresponds to one metric, with colors indicating the metric category: brown for complexity, blue for faithfulness, orange for localization, green for randomization, and red for robustness.
 Figure 12: Comparison of explanation methods based on evaluation metrics.
-(a): normalized average of all metrics across all RS datasets; (b): normalized average of the best-performing metric (CO, IROF, LLE, RRA, MPRT) for each category across all RS datasets.
+a): normalized average of all the considered metrics across all the considered RS datasets; b): normalized average of the best-performing metric (CO, IROF, LLE, RRA, MPRT) for each category across all considered RS datasets.
 Overall, GradCAM emerges as the most versatile explanation method, as it achieves strong performance across multiple evaluation categories.
 It excels in faithfulness, localization, and randomization, making it particularly well-suited for RS image scene classification.
-VI Conclusion and Guidelines
+V.C Evaluation of the computational efficiency
+TABLE V: Average computation time (in seconds) for metrics and explanation methods on the DG-ML dataset [46].
+Group Name Time (seconds)
+Explanation Methods DeepLIFT [12] 0.26
+GradCAM [13] 0.12
+LIME [10] 28.63
+LRP [11] 0.38
+Occlusion [9] 2.04
+Random baseline 0.04
+Explanation Metrics FE [28] 3.39
+IROF [29] 13.26
+AS [30] 12.15
+LLE [28] 11.97
+TKI [32] 0.10
+RRA [31] 0.16
+SP [33] 0.06
+CO [34] 0.06
+RL [36] 0.03
+MPRT [59] 11.10
+In this subsection, we analyze the computational efficiency of the considered metrics and explanation methods for RS scene classification.
+Table V reports the average runtime required in seconds to compute each explanation method or metric on the DG-ML dataset.
+From the table, one can observe that the average computation time required for the relevancy-propagation-based and gradient-propagation-based methods, such as GradCAM, DeepLIFT, and LRP, is below one second.
+This is because these methods explain the prediction in a single forward pass.
+In contrast, perturbation-based approaches such as LIME and Occlusion require substantially more computation, as they require multiple forward passes.
+The runtime of LIME is particularly high (28.63 seconds) due to the need to evaluate perturbed samples and fit a surrogate model.
+However, it is worth noting that the efficiency of perturbation-based methods depends mostly on their parametrization, as increasing the number of perturbations improves the reliability of the results but proportionally increases computation time.
+For explanation metrics, a similar behavior can be observed from the table.
+Perturbation-based metrics such as FE, IROF, AS, LLE, and MPRT require several seconds due to repeated recomputation of explanations under perturbations.
+In contrast, metrics from localization, complexity, and randomization categories, including TKI, RRA, SP, CO, and RL, are computed in closed form from a single attribution and thus achieve runtimes below 0.2 seconds.
+These results highlight that the computational efficiency of explanation metrics is largely dictated by whether perturbation is required.
+VI Conclusion and Discussion
 In this study, for the first time in RS, we have systematically investigated the suitability of explanation methods and metrics for RS image scene classification by combining methodological analysis with extensive experiments on three datasets.
@@
 Randomization metrics are often independent of the data.
-To support our methodological findings, we conducted two sets of experiments on three RS datasets: BEN-v2.0, DG-ML, and FBP-ML.
+To support our methodological findings, we conducted two sets of experiments on three RS datasets: BEN-v2.0 [60], DG-ML [46], and FBP-ML [62].
 In the first set of experiments, we meta-evaluated ten explanation metrics to assess their reliability.
-We observe that metrics from the robustness and randomization categories performed well.
-Conversely, as expected from the methodological analysis, metrics from the faithfulness and localization categories are more unreliable.
+We observe that metrics from the robustness and randomization categories are reliable.
+As expected from the methodological analysis, metrics from the faithfulness and localization categories are less reliable.
 In the second set of experiments, we use the most reliable explanation metrics to determine which explanation methods are best suited for RS image scene classification.
-Among the evaluated methods, GradCAM emerged as the most robust feature attribution method, demonstrating relatively stable performance across different datasets and evaluation metrics.
-However, it is important to note that no single method performed optimally across all evaluation criteria.
+Among the evaluated methods, GradCAM emerges as the most robust feature attribution method, demonstrating relatively stable performance across different datasets and evaluation metrics.
+However, it is important to note that no single method has performed optimally across all evaluation criteria.
 Following our methodological and experimental analysis, we develop guidelines for selecting suitable explanation metrics and explanation methods for scene classification in RS:
@@
 Complexity metrics should be interpreted carefully, as they depend strongly on the spatial extent of the target class.
-Similarly to localization metrics, complexity metrics could also be modified to account for the target class size.
+Similar to localization metrics, complexity metrics could also be modified to account for the target class size.
   4. 4.
-    We recommend using GradCAM as a general baseline due to its versatility and ease of interpretation.
+    We recommend using GradCAM as a general baseline due to its versatility, speed, and ease of interpretation.
 Among the evaluated methods, it consistently achieves high scores across multiple metric categories.
-As a final remark, we would like to note that, due to the lack of ground truth on explanations, it remains not possible to fully validate an explanation metric [24].
-The quantifiable reliability of a metric does not necessarily indicate its validity [66].
-Thus, each evaluation must be critically examined.
-Future work should focus on refining explanation metrics specifically designed for RS image scene classification.
-In particular, developing perturbation strategies that better accommodate spectral variability and feature redundancy may enhance the reliability of faithfulness metrics and, by extension, perturbation-based explanation methods.
-Although our study concentrated on feature attribution methods commonly used in RS (while being originally designed for natural images), a promising direction for future research is to develop and evaluate explanation techniques tailored specifically to the characteristics of RS images.
+Beyond methodological contributions, our findings have direct implications for operational applications.
+For instance, reliable explanation methods are crucial in maritime monitoring, where RS scene classification is used for maritime surveillance and decision support.
+Moreover, reliable attribution maps can serve as pseudo labels in tasks such as tree species classification [68].
+As a final remark, we emphasize that the evaluation protocol can not measure the intrinsic validity of explanation metrics, since no ground truth for explanations exists [24].
+It can probe the reliability and comparative behavior of metrics, yet reliability alone does not imply validity [67].
+Evaluation outcomes should therefore be regarded as indicative rather than definitive, and the theoretical foundations of the metrics also have to be taken into account.
+As a future work, we plan to refine explanation metrics specifically designed for RS image scene classification.
+In particular, we want to develop perturbation strategies that better accommodate spectral variability and feature redundancy, which may enhance the reliability of faithfulness metrics and, by extension, perturbation-based explanation methods.
+Although our study concentrated on feature attribution methods commonly used in RS (while being originally designed for natural images), a promising direction for future research is to develop explanation methods and metrics tailored specifically to the characteristics of RS images.
+Another important direction is to evaluate more recent explanation approaches, such as concept-based XAI, and to extend the evaluation framework accordingly.